### v1

# Latent Multimodal Reconstruction for Misinformation Detection

Abstract
Multimodal misinformation, such as miscaptioned images, where captions misrepresent an image’s origin, context, or meaning, poses a growing challenge in the digital age. To support fact-checkers, researchers have been focusing on creating datasets and developing methods for multimodal misinformation detection (MMD). Due to the scarcity of large-scale annotated MMD datasets, recent studies leverage synthetic training data via out-of-context image-caption pairs or named entity manipulations; altering names, dates, and locations. However, these approaches often produce simplistic misinformation that fails to reflect real-world complexity, limiting the robustness of detection models trained on them. Meanwhile, despite recent advancements, Large Vision-Language Models (LVLMs) remain underutilized for generating diverse, realistic synthetic training data for MMD. To address this gap, we introduce “MisCaption This!”, a training dataset comprising LVLM-generated miscaptioned images. Additionally, we introduce “Latent Multimodal Reconstruction” (LAMAR), a network trained to reconstruct the embeddings of truthful captions, providing a strong auxiliary signal to the detection process. To optimize LAMAR, we explore different training strategies (end-to-end training and large-scale pre-training) and integration approaches (direct, mask, gate, and attention). Extensive experiments show that models trained on “MisCaption This!” generalize better on real-world misinformation, while LAMAR sets new state-of-the-art on both NewsCLIPpings and VERITE benchmarks; highlighting the potential of LVLM-generated data and reconstruction-based approaches for advancing MMD. We release our code at: https://github.com/stevejpapad/miscaptioned-image-reconstruction.
Keywords Multimodal Learning  $\cdot$ Deep Learning  $\cdot$ Misinformation Detection  $\cdot$ Reconstruction Network
1 Introduction
The rise of the internet and digital technologies has significantly accelerated information dissemination while also amplifying the spread of misinformation, enabling new forms of deceptive content such as DeepFakes [1], multimodal misinformation [2], and LLM-generated misinformation [3]. Given the scale and rapid spread of misinformation, researchers have been developing automated fact-checking methods to support human fact-checkers in identifying false information more efficiently [4]. In this study, we focus on multimodal misinformation detection (MMD), specifically, the detection of misleading image-caption pairs, where texts and images jointly contribute to the spread of false or deceptive information [5].
Recent studies on MMD primarily focus on developing large-scale datasets and detection methods. Aside from a few small-scale annotated datasets [6, 7], existing datasets primarily consist of weakly annotated [8, 9, 10] or algorithmically generated datasets; created either by pairing images with out-of-context (OOC) captions from other images [11, 12, 13, 14] or by manipulating named entities to introduce inconsistencies, resulting in miscaptioned (MC) images [15, 16, 17]. On the modeling side, detection models have leveraged large pre-trained backbone encoders [14] or fine-tuned through self-supervised learning [18], improved modality fusion with attention-based techniques [19, 20, 21], incorporated external evidence [22, 23], and utilized Large Vision-Language Models (LVLMs) for both detection [24, 25] and explanation generation [26].
Nevertheless, named-entity manipulations often produce simplistic misinformation, lacking basic factual or logical consistency, while the potential of LVLMs to generate more diverse and robust synthetic training data remains unexplored. Moreover, despite the success of reconstruction networks in other domains, their potential remains largely unexplored for MMD where is particularly relevant, as the task closely aligns with human fact-checking practices, where reconstructing the true origin, context, or meaning of an image is a common strategy for debunking misinformation [27].
Figure 1: High-level overview of the proposed framework: an LVLM modifies a truthful image-caption pair, while LAMAR’s reconstruction network re-creates the original caption embedding produced by CLIP. This embedding is fused with other modalities using a mechanism (e.g., Gate) and fed into the detection network to produce the final verdict. The reconstruction network is trained to minimize the error between original and reconstructed embeddings, while the detection network is optimized for classification.
To this end, we propose a framework that leverages an LVLM to manipulate the original captions of images sourced from a dataset of truthful image-caption pairs, generating false captions that misrepresent aspects of the images. Subsequently, the proposed Latent Multimodal Reconstruction (LAMAR) network is tasked with reconstructing the embeddings of the original, truthful captions. As shown in Fig. 1, a truthful image-caption pair is manipulated by the LVLM, while LAMAR is tasked with reconstructing the original caption embedding using the image and the generated caption embeddings. Thereafter, the reconstructed embedding is integrated into the final detection network alongside the fused modalities. Our rationale is that LVLMs can generate more diverse and realistic training data, enhancing the generalizability of detection models, while the reconstruction process serves as an auxiliary signal, leveraging learned representations to refine embeddings by capturing consistency and inconsistency patterns in image-caption pairs—even in the absence of external evidence.
More specifically, we leverage an LVLM alongside “adversarial prompt selection” to limit the number of generative prompts by evaluating them against the LVLM’s zero-shot detection capabilities; filtering out prompts that produce easily detectable misinformation or overly generic image descriptions. This process results in the selection of four generative prompts, each used to generate a distinct version of our “MisCaption This!” dataset. For LAMAR, we employ a Transformer encoder with element-wise vector operations for modality fusion and explore two training strategies, end-to-end training and large-scale pre-training, as well as four methods for integrating the reconstructed embeddings into the detection network: direct integration, masking, gating, and self-attention.
Through extensive experiments with various training datasets, we demonstrate that models trained on “MisCaption This!” achieve superior out-of-distribution generalization on the VERITE evaluation benchmark [28], outperforming those trained on datasets relying on entity manipulation or cross-modal misalignment—with LAMAR performing 7.8% and 10.4% better, respectively. Moreover, LAMAR consistently outperforms prior state-of-the-art (SotA) models across training settings and datasets, achieving improvements of 4.3% on ‘True vs. MC’, 3.0% on ‘True vs. OOC’, and 4.8% on the multiclass task when trained on “MisCaption This!”.
2 Related Work
In recent years, automated fact-checking has attracted growing research interest [4], encompassing a range of challenging tasks such as claim detection [29], evidence filtering and retrieval [30, 31], fake news detection [32, 33, 34], retrieval of fact-checked articles [35, 36], DeepFake detection [1], and multimodal forms of misinformation [2]. In this paper, we focus on MMD, which is garnering increasing attention from researchers, leading to the ongoing development of datasets and methodologies for detecting false information and inconsistencies between images and their accompanying textual captions.
2.1 MMD Datasets
Training machine learning models for MMD requires suitable datasets, with current research focusing on annotated, weakly annotated, and synthetically generated data. Early MMD datasets, such as the ‘Twitter’ [7] and ‘Weibo’ [37] datasets, are relatively small and cover only a limited number of events-17 and 73, respectively-raising concerns about model generalization. To address this, larger weakly annotated datasets have emerged, including MuMiN [9], with rich social context but few images, and NewsBag [10], which includes satirical content, and Fakeddit [8], with over a million instances collected from Reddit. However, studies show that models trained on ‘Twitter’ and Fakeddit often exhibit unimodal biases, undermining their effectiveness in real-world multimodal misinformation detection [28].
Researchers have also explored synthetic data generation. These approaches can be categorized as either out-of-context (OOC) pairs or named entity swapping (NES). Early OOC datasets, like MAIM [11] and COSMOS [12], used random image-text mismatches, which often resulted in unrealistic and easy to detect samples [17]. More refined approaches, such as NewsCLIPings [14] and Twitter-COMMs [13], incorporated CLIP-based retrieval to enhance cross-modal relevance.
NES-based datasets generate misinformation by replacing named entities in captions with alternatives retrieved from similar or contextually relevant texts using cluster-based retrieval (MEIR [15]), rule-based substitutions (TamperedNews [16]), and CLIP-based retrieval (CLIP-NESt [17]). Despite advancements in LVLMs, only MMFakeBench leverages LVLM-generated rumors and AI-altered images and serves as a small evaluation benchmark [38]. Instead, we introduce a large LVLM-generated training dataset.
However, models trained and evaluated on synthetic data may struggle with real-world generalizability, as they may learn to detect patterns specific to artificially generated inconsistencies rather than the more complex and diverse manipulations found in real-world misinformation. To address this, benchmarks like VERITE incorporate real-world OOC and miscaptioned images [28].
2.2 MMD Methods
Research on MMD has centered on developing models that encode textual and visual modalities, fuse their representations, and assess their consistency and factual accuracy. Early methods, such as SpotFake [39], used VGG-19 and BERT, while recent approaches use pre-trained multimodal encoders such as CLIP [14], or fine-tune it through self-supervised learning [18]. Some models incorporate multi-task learning, such as EANN with an event discriminator [40] or MVAE, which uses an autoencoder to reconstruct the input text and visual features [41]; but does not modify the input text or reconstruct truthful captions from false ones.
While earlier methods relied on simple concatenation of visual and textual embeddings, more advanced approaches have explored Attention-based Multimodal Bilinear Pooling [19], Bidirectional Crossmodal Fusion (BCMF) [20], multi-head attention in Transformers [17], and element-wise vector fusion [21] to enhance cross-modal interaction. Recent work integrates external web evidence, with methods assessing internal and external consistency (CCN [22], SNIFFER [24]) or evaluating stance and relevance of external evidence (SEN [23], RED-DOT [21]). While external evidence is shown to improve performance, concerns remain about ‘leaked evidence’ from fact-checking articles [42, 31] and dataset artifacts that models may exploit instead of assessing factuality [43]. For these reasons, we do not consider evidence-based approaches in this study.
2.3 Reconstruction Networks
Reconstruction networks are deep learning models designed to generate or restore original data from low-resolution or altered inputs. They have been applied in various domains, including few-shot image classification by reframing it as a reconstruction problem in latent space [44], image inpainting through an adversarial framework guided by textual descriptions [45], super-resolution reconstruction to enhance low-resolution images of the same scene [46], and DeepFake detection [47]. Reconstruction networks remain largely unexplored in MMD, with MVAE [41] being a notable exception. However, MVAE uses an autoencoder to reconstruct the input text and image embeddings from a joint latent space but does not modify the text or generate truthful image descriptions from false ones; its objective is to extract more informative features for MMD within the latent space.
3 Problem Formulation
Given a set $\mathcal{D}^{t}={(I^{t}_{i},C^{t}_{i})}_{i=1}^{N}$ of $N$ image-caption pairs, where $I^{t}_{i}$ and $C^{t}_{i}$ represent an image and its matching, truthful caption, we define ‘Out-Of-Context’ (OOC) as any pair that combines an image $I^{t}_{i}$ with a caption $C^{x}_{i}$ taken from another image within $\mathcal{D}^{t}$; where x denotes the context. Similarly, we define ‘Mis-Captioned’ (MC) as any image whose accompanying caption $C^{f}_{i}$ has been manipulated to misrepresent the original, content, and/or meaning of the image; where “f” denotes falsehood.
We define a “Manipulator” as any method used to generate OOC pairs (e.g., CLIP-based retrieval) or MC pairs (e.g., named entities manipulation) from the original set $\mathcal{D}^{t}$ set, resulting in the sets $\mathcal{D}^{x}$ and $\mathcal{D}^{f}$, respectively.
We define MMD as a classification task with the objective to learn a mapping function $\mathsf{M}^{d}:\mathcal{D}\rightarrow\hat{y}$ where $\hat{y}$ represents the prediction upon target class $y$ in either of three scenarios:
1) Binary: ‘True vs. MC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{f}]$ with $K=N*2$ total pairs and $y\in\{0,1\}$
2) Binary: ‘True vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x}]$ with $K=N*2$ total pairs and $y\in\{0,2\}$,
3) Multi-class: ‘True vs. MC vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x},\mathcal{D}^{f}]$ with $K=N*3$ total pairs and $y\in\{0,1,2\}$.
Given image and text encoders $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$ producing image embeddings $\mathbf{I}$ and text embeddings $\mathbf{C}$ for image-caption pairs $(I,C)$ under examination, we define latent reconstruction as the task of learning a mapping function $\mathsf{M}^{r}:(\mathbf{I},\mathbf{C})\rightarrow\mathbf{\hat{C}^{t}}$, where $\mathbf{\hat{C}^{t}}$ represents the predicted embedding of the original, truthful caption $\mathbf{C^{t}}$. The model $\mathsf{M}^{r}$ is trained to minimize the reconstruction error between the true and predicted caption embeddings, such that:
$\mathcal{L}_{r}(C^{t},\hat{C}^{t})=\frac{1}{K}\sum_{i=1}^{K}(\mathbf{C}^{t}_{i}-\hat{\mathbf{C}}^{t}_{i})^{2}$ (1)
where $\mathcal{L}_{r}$ is the loss function (i.e., the Mean Squared Error (MSE)) that quantifies the discrepancy between the embeddings of the true and reconstructed caption embeddings.
Figure 2: Examples of truthful and generated captions taken from the “MisCaption This!” ($\mathcal{D}_{3}$) dataset, alongside false captions created via named entity swaps.
4 Construction of “MisCaption This!”
In this study, we explore the creation of a synthetic training dataset of miscaptioned images ($\mathcal{D}^{f}$) by manipulating the image captions of a truthful dataset ($\mathcal{D}^{t}$) using an LVLM as the “Manipulator”. Our rationale is that LVLMs, with their advanced multimodal understanding and generation capabilities, can produce more realistic false captions for images compared to methods relying on manipulating named entities. In turn, we hypothesize that the generated data $\mathcal{D}^{f}$ can be leveraged to train more robust detection models $\mathsf{M}^{d}$, thus enhancing their ability to generalize to real-world misinformation.
4.1 Generative Model
To generate synthetic data, we employ LLaVa-1.6 (Large Language and Vision Assistant) [48] from Hugging Face [^1] https://huggingface.co/llava-hf/llava-v1.6-mistral-7b-hf leveraging Mistral-7B-Instruct-v0.2. We also explored LLaVa-1.6 (Vicuna-13B), Janus Pro 7B (DeepSeek), and MiniGPT-v2 (Llama-2-7B) as alternatives, however, we were unable to get them to consistently generate realistic false captions, as they often defaulted to generic image captioning, re-phrasing of the original caption, or overly simplistic misinformation. In contrast, GPT-4o Mini demonstrated robust safeguards, often refraining from generating misinformation altogether. Therefore, this part of our work serves as an initial exploration of LVLM-generated synthetic data, with future research needed to assess the impact of other LVLMs and prompting methods.
4.2 Adversarial Prompt Selection
Generating realistic misinformation with an LVLM requires a carefully selected generative prompt $p^{gen}$. However, determining which $p^{gen}$ will produce a dataset $\mathcal{D}^{f}$ that leads to better-performing detection models $\mathsf{M}^{d}$ is not straightforward and would require significant resources for an extensive empirical examination. To address this, we propose Adversarial Prompt Selection, which refines the set of generative prompts by assessing their generated captions $C^{f}$ based on their ability to evade detection by a zero-shot LVLM, ensuring that the generated data presents a meaningful challenge for detection models, rather than being trivially misleading or overly simplistic.
First, we evaluate the LVLM’s ability to perform zero-shot detection of miscaptioned images. We experiment with multiple detection prompts for LLaVa, We experiment with multiple detection prompts for LLaVa, varying their formulation (e.g., word synonyms, rephrasing) and compared their zero-shot performance on a subset of 100 real-world “True vs. MC” samples from VERITE. Based on this process, the detection prompt ($p^{dt}$): “The image is captioned as: [CAPTION]. Is the caption truthful or does it contain falsehoods?” was selected for the next stage.
Next, we evaluate a range of generative prompts against zero-shot LLaVa with $p^{dt}$, experimenting with various formulations similar to those used for the detection prompts. For each generative prompt ($\forall j\in J$), LLaVa generates 1,000 captions from a randomly selected subset of 1,000 truthful image-caption pairs from the NewsCLIPpings training set. Then, the LVLM is tasked with classifying both the truthful and generated captions, and we compute the average detection accuracy for each prompt. This process can be formulated as:
$C^{f}_{j}=\text{LLaVa}(I,C^{t}\mid p^{gen}_{j}),\quad\forall j\in[1,J]$ (2)
$\hat{y}_{j}=\text{LLaVa}(I,C_{j}\mid p^{dt}),\quad C_{j}\in\{C^{t}_{j},C^{f}_{j}\}$ (3)
We reason that generated captions produced by simplistic prompts, which are easily detected, indicate the generation of unrealistic misinformation and are unlikely to contribute to training effective $\mathsf{M}^{d}$ models. For instance, generated captions from prompts such as ‘Create a false caption for the image to propagate misinformation’ and ‘Create fake news based on the image’ were detected with 83.5% and 81.0% accuracy, respectively. Conversely, generative prompts that produce undetectable captions, may fail to produce actual misinformation, instead leaving the original captions unchanged or rephrased. For example, the prompts ‘Provide a caption for the image’ and ‘Maintain the original caption of the image’ resulted in low detection accuracy (42.5% and 50.0%) while having low (7%) or perfect (100%) word overlap with the original captions, indicating a generic or truthful image description rather than misinformation.
Through this process, and after manual inspection, we select four generative prompts, $p^{gen}_{1}$, $p^{gen}_{2}$, $p^{gen}_{3}$, and $p^{gen}_{4}$, with varying adversarial zero-shot detection performance: 57%, 62%, 71%, and 75%. These prompts are then used to generate four distinct sets, $\mathcal{D}^{f}_{1}$, $\mathcal{D}^{f}_{2}$, $\mathcal{D}^{f}_{3}$, and $\mathcal{D}^{f}_{4}$, each corresponding to a different set of false captions related to the images of $\mathcal{D}^{t}$. All four prompts are multimodal, as they take into account both the image and its original caption to generate false captions. For ethical considerations, we refrain from disclosing the specific prompts here, as sharing them could contribute to the automated generation of misinformation; however, they will be available upon request under a research-only license.
4.3 Generated Data Filtering
Through manual examination, we observe that generative prompts produce more coherent and plausible false captions, often better than NES-based methods in terms of logical coherence and knowledge-based consistency. For instance, as shown in Fig.2, manipulating named entities led to noticeable inconsistencies, such as the incorrect claims of the “Indian parliament” being in “Bangladesh” and “Skeldergate” being in “San Francisco”. These examples underscore the limitations of NES-based methods in preserving basic logical and factual consistency. In contrast, the LVLM demonstrated the ability to generate more creative and plausible misinformation. For instance, it misrepresented a demonstration by reframing its purpose, shifting it from anti-austerity to pro-austerity, and further reinforced the false narrative with claims of “successful implementation” and “widespread approval”. Similarly, it distorted the aftermath of the evacuation following heavy flooding, suggesting that people were “enjoying a leisurely swim in the flooded streets” and “turning a disaster into a fun-filled day”.
However, we also observed that LLaVa can occasionally “ramble”, often rephrasing the original caption while adding redundant or superficial details. For instance, given the original caption “The recent economic boom has enabled new projects such as the Union Trade Centre shopping centre in the heart of Kigali”, LLaVa generated: “The Union Trade Centre shopping centre in Kigali is a prime example of the city’s thriving economy with numerous cars parked outside indicating a bustling shopping scene”; thereby rephrasing the original caption and incorporating minor descriptive elements that do not amount to misinformation.
To address this issue, we apply a post-processing filter that removes samples where the length of the generated caption exceeds a specified character threshold $l\in{\{0,5,10,15,25,50,None\}}$ relative to the original, truthful caption. To maintain a balanced dataset, we remove both the generated pairs ($I^{t}_{i},C^{f}_{i}$) and their corresponding truthful pairs ($I^{t}_{i},C^{t}_{i}$) that exceed this threshold. This filtering process retains 4.5%, 19.1%, 27.8%, 34.9%, 47.9%, 74.0%, and 100% of the dataset for $l\in{\{0,5,10,15,25,50,None\}}$, respectively. We empirically evaluate the impact of this filtering on model performance.
4.4 Dataset Source and Statistics
We use the NewsCLIPpings ‘Merged/Balanced’ version [14], which has been shown to be effective for OOC detection [22, 24, 21, 43], as the source dataset $\mathcal{D}^{t}$ from which to generate $\mathcal{D}^{f}$, utilizing only its truthful data: 35,536 pairs for training, 3,512 for validation, and 3,512 for testing. After generating $\mathcal{D}^{f}_{1}$, $\mathcal{D}^{f}_{2}$, $\mathcal{D}^{f}_{3}$, and $\mathcal{D}^{f}_{4}$ from $\mathcal{D}^{t}$ using LLaVa and the corresponding generative prompts, we merge $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{f}]$ to address the “True vs. MC” task, or integrate the full NewsCLIPpings dataset $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x},\mathcal{D}^{f}]$ to address the multi-class task. We preserve the original train/validation/test split of the NewsCLIPpings dataset to prevent any data leakage. This results in a total of 106,605, 10,536, and 10,896 samples for training, validation, and testing, respectively, ensuring a balanced distribution across the three classes.
5 Latent Multimodal Reconstruction (LAMAR)
Our objective is to develop a reconstruction network $\mathsf{M}^{r}$ that utilizes the embeddings $\mathbf{I},\mathbf{C}$ of image-caption pairs to reconstruct the embedding of the original, truthful caption associated with the image. The reconstructed embeddings $\mathbf{\hat{C}^{t}}$ are then integrated within the final detection model, along with the fused modalities, to aid the detection process. Our rationale is that if the reconstructed embedding $\mathbf{\hat{C}^{t}}$ closely match the input embeddings $\mathbf{C}$, then, the pair $\mathbf{I},\mathbf{C}$ is likely to be truthful. On the other hand, significant discrepancies between $\mathbf{C}$ and $\mathbf{\hat{C}^{t}}$ should suggest potential manipulation of $C$. We hypothesize that this process will provide valuable signals for the detection model.
Since reconstruction networks remain under-explored in the context of MMD, we investigate various alternative strategies, broadly categorized into: (1) end-to-end training, and (2) large-scale pre-training. Additionally, we explore attention, gating, and masking mechanisms to refine the integration of the reconstructed embedding into the Detection network. As illustrated in Figure 3, our pipeline consists of a backbone encoder, followed by modality fusion, a transformer-based reconstruction module, integration of the reconstructed embedding, and a final classification layer. We provide a detailed discussion of these components in the following sections.
Figure 3: End-to-end training of the proposed LAMAR architecture, which utilizes a CLIP ViT L/14 backbone encoder and a Transformer encoder for the reconstruction network with element-wise vector operations for enhanced modality fusion. The reconstructed embedding is then integrated into the detection network through various mechanisms (e.g., gate, mask, attention, or no mechanism) which predicts the final verdict. The reconstruction network is optimized using MSE loss to minimize the difference between reconstructed and ground truth caption embeddings, while the detection network is optimized with cross-entropy (CE) loss.
5.1 Backbone Encoder
We use CLIP ViT L/14 from OpenCLIP  [^2] https://github.com/mlfoundations/open_clip as the backbone multimodal encoder, $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$, to produce image embeddings $\mathbf{I}\in{R}^{768\times 1}$ and text embeddings $\mathbf{C}\in{R}^{768\times 1}$, which are pre-aligned within a shared embedding space.
5.2 Modality Fusion
To effectively integrate visual and textual modalities, we employ a fusion strategy that combines concatenation (;) with element-wise vector operations (addition, subtraction, and multiplication). This approach has been shown to be a lightweight yet effective method for capturing complementary relationships and differences between the two modalities [21]. Specifically, we define the fused representation $\mathbf{F}$:
$\mathbf{F}=[\mathbf{I};\mathbf{I}+\mathbf{C};\mathbf{I}-\mathbf{C};\mathbf{I}*\mathbf{C};\mathbf{C}]$ (4)
with $\mathbf{F}\in{R}^{5\times 768}$.
5.3 Detection Network
For the detection model $\mathsf{M}^{d}$, we define a neural network as follows:
$\hat{y}=\textbf{W}_{1}\cdot\text{GELU}(\textbf{W}_{0}\cdot\text{Flatten}([\mathbf{F};\mathbf{\hat{C}^{t}}]))$ (5)
where $\textbf{W}_{0}\in\mathbb{R}^{768\times 768}$ is a fully connected layer followed by a GELU activation, and $\textbf{W}_{1}\in\mathbb{R}^{n\times 768}$ is the final classification layer, with $n=1$ for binary classification or $n=3$ multi-class classification. Bias terms $b$ are included in the model but omitted here for brevity. Here, the operation ‘Flatten’ refers to converting the concatenated vector $[\mathbf{F};\mathbf{\hat{C}^{t}}]$ into a one-dimensional vector before passing it through the hidden layers. The Detection Network is optimized using binary cross-entropy loss for binary classification and categorical cross-entropy for multi-class classification; denoted as $\mathcal{L}_{d}$.
5.4 Reconstruction Network
For the reconstruction network $\mathsf{M}^{r}$, we follow prior research in using Transformer encoder $\mathsf{T(\cdot)}$ for MMD [17, 28, 21, 43], formulated as:
$[\mathbf{t}_{\text{CLS}},\mathbf{t}_{F}]=\mathsf{T}([\mathbf{CLS};\mathbf{F}]),\quad\mathbf{\hat{C}^{t}}=\mathbf{t}_{\text{CLS}}$ (6)
where $CLS$ is a trainable classification token that serves as a global representation of all inputs, and its transformation is defined as the reconstructed caption embedding $\mathbf{\hat{C}^{t}}$. The network is optimized using the MSE loss function, as defined in Eq. 1.
5.4.1 End-to-end (E2E) training
To jointly optimize the reconstruction network $\mathsf{M}^{r}$ and the detection model $\mathsf{M}^{d}$, we explore end-to-end multi-task training, where the entire LAMAR model is trained using a joint loss function, which combines the reconstruction objective $\mathcal{L}_{r}$ (e.g., MSE) and the detection loss $\mathcal{L}_{d}$ (binary or categorical cross-entropy) as $\mathcal{L}=\mathcal{L}_{d}+\mathcal{L}_{r}$.
This approach enables the detector to utilize both the input representation and the reconstructed embedding, while we explore various mechanisms to dynamically adjust the contribution of the reconstructed embedding based on relevance. Specifically, we explore:
1.1) Direct integration of the reconstructed caption embedding $\mathbf{\hat{C}^{t}}$ into $\mathsf{M}^{d}$, as shown in Eq.5.
1.2) A gating mechanism, formulated as:
$\mathbf{g}=\mathcal{S}(\mathbf{W}_{g}\cdot\mathbf{F}+\mathbf{b}_{g})$ (7)
$\mathbf{\hat{C}^{t}}_{gate}=\mathbf{g}\odot\mathbf{\hat{C}^{t}}$ (8)
where $\mathbf{g}$ is the gate, $S$ the sigmoid function and $\mathbf{W_{g}}\in\mathbb{R}^{768\times 768}$ while the input to Eq.5 is altered to $[\mathbf{F};\mathbf{\hat{C}^{t}}_{gate}]$.
1.3) A masking mechanism, formulated as:
$\mathbf{m}\sim Bernoulli(\mathcal{S}(\mathbf{W}_{m}\cdot\mathbf{F}+\mathbf{b}_{m}))$ (9)
$\mathbf{\hat{C}^{t}}_{mask}=\mathbf{m}\odot\mathbf{\hat{C}^{t}}$ (10)
where $\mathbf{m}$ represents binary mask sampled from a Bernoulli distribution, $\mathbf{W_{m}}\in\mathbb{R}^{768\times 768}$ and the input to Eq.5 is altered to $[\mathbf{F};\mathbf{\hat{C}^{t}}_{mask}]$.
1.4) An attention mechanism, formulated as:
$\mathbf{F_{a}}=[\mathbf{I},\mathbf{C},\mathbf{\hat{C}}]$ (11)
$\mathbf{Q}=\mathsf{W}_{Q}\cdot\mathbf{F_{a}},\quad\mathbf{K}=\mathsf{W}_{K}\cdot\mathbf{F_{a}},\quad\mathbf{V}=\mathsf{W}_{V}\cdot\mathbf{F_{a}}$ (12)
$\mathbf{\hat{C}^{t}}_{attend}=mean(softmax\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{T}}{\sqrt{768}}\right)\cdot\mathbf{V})$ (13)
where $\mathsf{W}_{Q},\mathsf{W}_{K},\mathsf{W}_{V}\in\mathbb{R}^{768\times 768}$, and $mean$ denotes average pooling across the first dimension. The input to Eq.5 is altered to $[\mathbf{F};\mathbf{\hat{C}^{t}}_{attend}]$.
5.4.2 Large-scale Pre-Training (PT)
In addition to end-to-end training, we also investigate a large-scale pre-training approach where the reconstruction network is trained exclusively on truthful captions using a large-scale image-caption dataset, VisualNews [49], comprising 1,259,732 truthful image-caption pairs. We explore two pre-training strategies:
2.1) Gaussian noise is added to the original text embedding, and the network is tasked with reconstructing the noisy embedding; expressed as:
$\mathbf{C}^{f}=\mathbf{C}^{t}+\mathcal{N}(\mu,\sigma^{2})$ (14)
where $\mathcal{N}(\mu,\sigma^{2})$ denotes Gaussian noise with mean $\mu$ and standard deviation $\sigma$.
2.2) Similarly, dropout is applied to the original text embedding, and the network is tasked with reconstructing the dropped-out embedding; expressed as:
$\mathbf{C}^{f}=Dropout(\mathbf{C}^{t},dp)$ (15)
where $dp$ denotes the dropout probability.
In both cases, the noisy or dropout-modified embedding $\mathbf{C^{f}}$ is then substituted for $\mathbf{C}$ in Eq. 4 and used in $\mathsf{M}^{r}$ to reconstruct the truthful caption embedding $\mathbf{\hat{C}}^{t}$. Once $\mathsf{M}^{r}$ is trained, the reconstructed embeddings are integrated into the detection network $\mathsf{M}^{d}$ during its training. To integrate these embeddings, we explore direct integration as well as the gate and attention mechanisms.
6 Experimental Setup
6.1 Training Datasets
For the “True vs. MC” experiments, we use the four versions of our “MisCaption This!” dataset, as detailed in Section 4, along with two additional datasets: the deduplicated version of the Crossmodal HArd Synthetic MisAlignment (CHASMA) dataset, which includes 145,891 truthful and 145,891 miscaptioned images [28], and the CLIP-based Named Entity Swapping by Topic (NESt) dataset, containing 847,693 miscaptioned images and 1,007,744 truthful images [17]. For “True vs. OOC” experiments, we use the NewsCLIPpings dataset [14], Merged/Balanced version, comprising 42,680 truthful and OOC samples in total. Finally, for multi-class classification, we combine one of the “miscaptioned images” datasets (“MisCaption This!”, CHASMA, or NESt) with the NewsCLIPpings dataset, which represents the OOC class, and apply random under-sampling to balance the classes.
6.2 Evaluation Protocol
We adhere to the training, validation, and testing splits provided by each dataset and evaluate models using Accuracy as the primary metric. After training on any of the aforementioned synthetic datasets, we further assess performance on the VERITE benchmark [28], which comprises 1,000 real-world samples: 338 truthful image-caption pairs, 338 miscaptioned images, and 324 out-of-context pairs. For binary classification, we report “True vs. OOC” and “True vs. MC” accuracies, while for the multi-class task, we provide overall accuracy along with per-class accuracy. All three tasks serve as out-of-distribution evaluations, as the training data are synthetically generated, while the final evaluation is conducted on real-world data.
6.3 Competing Methods
In addition to our proposed LAMAR architecture and its variations, we reproduce and evaluate the performance of several competing methods, without using any external evidence: (1) DT-Transformer [17]: a Transformer encoder that processes the input sequence $[\mathbf{CLS};\mathbf{I};\mathbf{C}]$; (2) RED-DOT (‘Baseline’ version) [21]: a Transformer-based model utilizing the fused representation $[\mathbf{CLS};\mathbf{F}]$ as input; (3) AITR (Attentive Intermediate Transformer Representations) [43], a model incorporating multimodal similarity (MUSE) and a self-attention mechanism over a stack of Transformer encoder blocks, with varying numbers of multi-head self-attention layers. We consider both its “attention pooling” and “weighted pooling” variants.
6.4 Implementation Details
We train LAMAR using a Transformer encoder with 4 encoder blocks, each containing 4 multi-head self-attention heads and a feed-forward layer of 1,024 dimensions and a dropout of 0.1 probability. The model is optimized with Adam, using a learning rate of $41e-$ 4 and a batch size of 512. We train the network for up to 50 epochs, with early stopping after 10 epochs if validation performance does not improve. For the large-scale pre-training, we consider $\sigma\in\{0.1,0.2\}$ and $\mu=0.0$ and dropout probability $dp\in\{0.2,0.5\}$. To ensure reproducibility, we set a constant random seed of 0 for PyTorch, Python Random, and NumPy.
6.5 Computational Complexity
Generating a single version of “MisCaption This!”  including the generation of captions for 35,536 images, took 18.3 hours using a single Nvidia GeForce RTX 4090 (24GB RAM). Feature extraction with CLIP ViT-L/14 required approximately 16 minutes for images and 2 minutes for texts, using a batch size of 256. Both data generation and feature extraction were repeated four times for each version of “MisCaption This!”.
Using fvcore [^3] https://github.com/facebookresearch/fvcore, we estimate that LAMAR, trained end-to-end with a gate mechanism, requires 980,997 FLOPs, and 1,034,181 FLOPs with the attention mechanism. This is comparable to AITR (1,070,653 FLOPs) [43] and significantly more efficient than RED-DOT (2,208,072 FLOPs) [21]. When using pre-extracted features, LAMAR requires a maximum of 30 seconds per epoch (batch size 512) to process the full “MisCaption This!” for multiclass classification.
Figure 4: Performance of detection models (DT-Transformer and RED-DOT) trained on four variations of the “MisCaption This!” dataset ($\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\mathcal{D}_{3}$, and $\mathcal{D}_{4}$), evaluated under varying filtering thresholds ($l\in{0,5,10,15,25,50,\text{None}}$) in terms of test-set accuracy and VERITE “True vs. MC” accuracy.
7 Results
7.1 Dataset Variants and Filtering
We first examine the generalizability of two established MMD models: DT-Transformer [17] and RED-DOT [21] (without incorporating external evidence) when trained on the four “MisCaption This!” variations ($\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\mathcal{D}_{3}$, and $\mathcal{D}_{4}$), under different filtering thresholds $l\in\{0,5,10,15,25,50,\text{None}\}$. As shown in Fig. 4, test-set accuracy tends to improve with higher $l$ values across all datasets and models. The highest performance is observed with no filtering ($l=\text{None}$), ranging from 79% to 83.4%. In contrast, evaluating models on the real-world data of VERITE (“True vs. MC”), there is a slight negative (Pearson) correlation between test-set accuracy and VERITE accuracy: -0.29 for DT-Transformer and -0.17 for RED-DOT. This suggests that higher test-set accuracy does not guarantee better generalization, as training with synthetic misinformation that is too easily distinguishable from truthful pairs may lead to inflated test-set accuracy without improving performance on real-world misinformation.
Both models achieve peak performance at $l=10$, with accuracies of 63.3% and 61.8%, followed by $l=5,15,$ and $25$; highlighting the importance of filtering out “rambling” and generic descriptions sometimes produced by LLaVa. Thus, by refining the dataset in this way, we improve model robustness and enhance real-world generalization on VERITE. Overall, both models achieve their highest performance when trained on $\mathcal{D}_{3}$, with RED-DOT and DT-Transformer averaging 60.2% and 59.2%, respectively. This suggests that $p_{3}^{gen}$ generates misinformation patterns that more closely resemble real-world examples, leading to better model generalization. Therefore, for subsequent experiments, we will continue using the $\mathcal{D}_{3}$ version of “MisCaption This!”, while treating $l$ as a tunable hyper-parameter.
Table 1: Comparative and ablation analysis of the proposed LAMAR method and prior SotA models (DT-Transformer, RED-DOT, and AITR) trained on three synthetic datasets (“MisCaption This!” ($\mathcal{D}_{3}$), NESt, and CHASMA) and evaluated on VERITE (“True vs. MC”). Bold denotes the highest accuracy on VERITE.
Model $\mathcal{D}_{3}$ NESt CHASMA
DT-Transformer 61.8 58.7 57.8
RED-DOT 63.3 57.0 58.0
AITR 62.9 58.4 58.3
LAMAR Variant $\mathcal{D}_{3}$ NESt CHASMA
E2E, Gate 66.0 61.2 59.8
E2E, Attention 65.1 58.6 58.6
E2E, Mask 63.6 59.5 58.1
E2E, Direct 63.8 59.2 57.8
E2E, Direct-Image 63.2 58.3 57.3
E2E, Direct-Text 63.0 58.7 57.1
No Reconstruction 62.7 58.2 56.7
PT, Dropout, Gate 63.0 57.5 58.6
PT, Dropout, Attention 64.5 57.1 56.5
PT, Dropout, Direct 61.2 59.5 57.1
PT, Gaussian, Gate 62.1 58.7 59.3
PT, Gaussian, Attention 63.0 59.3 57.5
PT, Gaussian, Direct 63.2 58.6 57.5
Table 2: Performance of models on the “True vs. OOC” task, trained on the NewsCLIPpings dataset and evaluated on both NewsCLIPpings and VERITE.
Model NewsCLIPpings VERITE
CLIP [14] 60.2 -
SSDL [18] 71.0 -
DT-Transformer 79.7 69.4
RED-DOT 81.5 73.5
AITR 84.1 74.1
LAMAR [E2E, Attention] 84.8 75.1
LAMAR [E2E, Gate] 84.7 76.3
7.2 Ablation across Datasets
We train three established MMD models, DT-Transformer, RED-DOT, and AITR, alongside various variants and ablations of the proposed LAMAR method using three synthetic “True vs. MC” training datasets: “MisCaption This!” ($\mathcal{D}_{3}$), CHASMA [28], and NESt [17], and evaluating their performance on VERITE (“True vs. MC”).
As shown in Table 1, using “MisCaption This!” ($\mathcal{D}_{3}$) as the training dataset consistently results in models with higher generalizability to real-world data (VERITE). This trend holds across all evaluated models and variants. These results support our hypothesis that LVLMs are capable of generating more effective synthetic training data, ultimately leading to more robust detection models, compared to previous approaches that rely on named entity swaps or cross-modal misalignment. Furthermore, we observe that LAMAR trained end-to-end (E2E) with the “Gate” mechanism for integration achieves the highest overall performance on VERITE across all three datasets, yielding the best performance (66%) when using the “MisCaption This!” ($\mathcal{D}_{3}$); achieving 7.8% higher performance compared to when trained on NESt, and 10.4% higher compared to CHASMA. Overall, we observe that E2E training tends to yield better performance, as it is directly optimized to reconstruct the embeddings of actual miscaptioned images, in contrast to PT methods that reconstruct embeddings of truthful pairs interjected with Gaussian noise or dropout.
Additionally, LAMAR with E2E training, utilizing either the “Gate” or “Attention” mechanisms for integrating the reconstruction embeddings, consistently outperforms the direct integration; emphasizing the importance of utilizing an integration mechanism. The “Mask” mechanism does not perform as well, likely due to the complete masking of the reconstruction embedding, which reduces its expressiveness. Instead, modulating the embedding by adjusting its values based on learned gate values or attention scores proves to be a more effective strategies.
We also investigate unimodal reconstruction as an ablation. Direct integration without the image input (Direct-Image), using only the text as input, results in a slight decrease in performance (-0.9%). Similarly, excluding textual information (Direct-Text), which effectively transforms the task into a latent image captioning approach, leads to even lower performance (-1.3%). Removing the reconstruction network entirely from LAMAR, leaving only the detection classifier, yields the lowest performance among all end-to-end (E2E) approaches. Finally, we observe that LAMAR [E2E, Gate] outperforms prior state-of-the-art models, such as DT-Transformer, RED-DOT, and AITR, by 6.8%, 4.3%, and 4.9%, respectively, when trained on “MisCaption This!” ($\mathcal{D}_{3}$); validating the effectiveness of our proposed method for the detection of miscaptioned images.
Table 3: Performance of models trained on the multi-class “MisCaption This!” ($\mathcal{D}_{3}$) or the combined CHASMA and NewsCLIPpings datasets, evaluated on VERITE, reported as overall accuracy and per-class accuracy.
Training Model Accuracy True MC OOC
CHASMA DT-Transformer 50.0 78.7 23.1 48.0
RED-DOT 48.5 71.9 21.9 51.7
AITR 51.4 90.5 15.4 48.0
LAMAR [E2E, Gate] 53.2 89.6 18.1 51.7
$\mathcal{D}_{3}$ DT-Transformer 47.8 53.0 60.4 29.2
RED-DOT 48.8 50.9 53.3 41.9
AITR 51.7 50.3 62.7 41.5
LAMAR [E2E, Gate] 54.2 58.6 58.6 44.9
7.3 Out-of-context and Multiclass Detection
Table 2 presents the performance of LAMAR compared to prior SotA models on the binary classification task of “True vs. OOC”, without external evidence. LAMAR achieves the highest accuracy on the NewsCLIPpings dataset, with LAMAR [E2E, Attention] reaching 84.8% and LAMAR [E2E, Gate] closely following at 84.7%, outperforming AITR (84.1%) and other models. More importantly, when evaluated on VERITE, LAMAR [E2E, Gate] attains 76.3%, surpassing the best prior model (AITR) by 3%. We observe a noticeable performance gap between “True vs. OOC" (76.3%) and “True vs. MC" (66.0%) due to differing task complexity. In OOC, the entire image is mismatched with the caption, making discrepancies more apparent, while MC involves subtle manipulations (e.g., actions, dates, locations, names), making it harder to distinguish from truthful captions.
For the multi-class classification task, we consider two training datasets: “MisCaption This!” ($\mathcal{D}_{3}$) and CHASMA combined with NewsCLIPpings. As shown in Table 3, LAMAR [E2E, Gate] continues to demonstrate superior generalization, regardless of the dataset used for training. When trained on “MisCaption This!” ($\mathcal{D}_{3}$), LAMAR achieves the highest overall accuracy (54.2%), representing a relative improvement of 4.8%, 11.1%, and 13.4% over AITR, RED-DOT, and DT-Transformer, respectively. Notably, all four models achieve relatively higher accuracy on the MC class when trained on “MisCaption This!” ($\mathcal{D}_{3}$) compared to CHASMA+NewsCLIPpings. These results validate the effectiveness of our proposed method in leveraging synthetic data generated by an LVLM and highlight the impact of incorporating a reconstruction network to enhance the detection process, ultimately improving generalization to real-world misinformation detection.
8 Conclusions
In this study, we address the challenge of MMD by introducing the “MisCaption This!” dataset, and the Latent Multimodal Reconstruction (LAMAR) method. To create “MisCaption This!”, we use an LVLM to generate diverse and realistic synthetic miscaptioned images, enhancing training data quality and by extension model generalization. LAMAR employs a reconstruction-based approach, where the original truthful caption embeddings are reconstructed from manipulated image-caption pairs, providing an auxiliary signal to aid the detection process. Extensive experiments show that models trained on “MisCaption This!” achieve superior out-of-distribution generalization on the real-world VERITE benchmark, outperforming named entity manipulations and cross-modal misalignment by 7.8% and 10.4%, respectively. Additionally, LAMAR set new state-of-the-art performance on all three VERITE tasks—“True vs. OOC”, “True vs. MC”, and multiclass classification—outperforming prior SotA by 4.3%, 3.0%, and 4.8%, respectively. These results highlight the potential usefulness of leveraging LVLMs for dataset generation and the potential of reconstruction networks in improving MMD performance.
Our study represents a first step toward leveraging LVLMs to enhance training data quality for MMD but is limited to leveraging a single model, LLaVa 1.6 7B. Future research could explore alternative LVLMs, jailbreak prompts, prompt tuning or few-shot prompting to further improve the robustness of generated training data. Nevertheless, researchers should consider potential risks, particularly in inadvertently enabling malicious actors to refine AI-generated disinformation campaigns. Furthermore, future research could further refine the reconstruction network by integrating additional external information or evidence to enhance the detection process. We do not explore this avenue due to the absence of datasets containing external evidence for miscaptioned images. A key challenge in collecting and leveraging external information is the risk of information leakage, which can undermine the early detection of emerging misinformation [42, 31]. Additionally, existing evidence-based OOC datasets often contain artifacts that models may exploit as shortcuts rather than truly assessing factual consistency [43]. Addressing these limitations would be a valuable step forward for the fields of MMD and automated fact-checking.
Acknowledgments
This work is partially funded by the project “vera.ai: VERification Assisted by Artificial Intelligence” under grant agreement no. 101070093 and “DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies” under grant agreement no. 101079164.

### v3

@@
-Latent Multimodal Reconstruction for Misinformation Detection
+Latent Reconstruction from Generated Data for Multimodal Misinformation Detection
 Abstract
 Multimodal misinformation, such as miscaptioned images, where captions misrepresent an image’s origin, context, or meaning, poses a growing challenge in the digital age.
-To support fact-checkers, researchers have been focusing on creating datasets and developing methods for multimodal misinformation detection (MMD).
-Due to the scarcity of large-scale annotated MMD datasets, recent studies leverage synthetic training data via out-of-context image-caption pairs or named entity manipulations; altering names, dates, and locations.
-However, these approaches often produce simplistic misinformation that fails to reflect real-world complexity, limiting the robustness of detection models trained on them.
-Meanwhile, despite recent advancements, Large Vision-Language Models (LVLMs) remain underutilized for generating diverse, realistic synthetic training data for MMD.
-To address this gap, we introduce “MisCaption This!”, a training dataset comprising LVLM-generated miscaptioned images.
-Additionally, we introduce “Latent Multimodal Reconstruction” (LAMAR), a network trained to reconstruct the embeddings of truthful captions, providing a strong auxiliary signal to the detection process.
-To optimize LAMAR, we explore different training strategies (end-to-end training and large-scale pre-training) and integration approaches (direct, mask, gate, and attention).
-Extensive experiments show that models trained on “MisCaption This!” generalize better on real-world misinformation, while LAMAR sets new state-of-the-art on both NewsCLIPpings and VERITE benchmarks; highlighting the potential of LVLM-generated data and reconstruction-based approaches for advancing MMD.
-We release our code at: https://github.com/stevejpapad/miscaptioned-image-reconstruction.
+Due to the scarcity of large-scale annotated datasets for multimodal misinformation detection (MMD), recent approaches rely on synthetic training data created via out-of-context pairings or named entity manipulations (e.g., altering names, dates, or locations).
+However, these often yield simplistic, unrealistic examples, which limits their utility as training examples.
+To address this, we introduce “MisCaption This!”, a framework for generating high-fidelity synthetic miscaptioned datasets through Adversarial Prompting of Vision-Language Models (VLMs).
+Additionally, we introduce “Latent Multimodal Reconstruction” (LAMAR), a Transformer-based network trained to reconstruct the embeddings of truthful captions, providing a strong auxiliary signal to guide detection.
+We explore various training strategies (end-to-end vs.
+large-scale pre-training) and integration mechanisms (direct, mask, gate, and attention).
+Extensive experiments show that models trained on “MisCaption This!” data generalize better to real-world misinformation, while LAMAR achieves new state-of-the-art on NewsCLIPpings, VERITE, and the newly introduced VERITE 24/25 benchmark; highlighting the efficacy of VLM-generated data and reconstruction-based networks for advancing MMD.
+Our code is available at https://github.com/stevejpapad/miscaptioned-image-reconstruction.
 Keywords Multimodal Learning $\cdot$ Deep Learning $\cdot$ Misinformation Detection $\cdot$ Reconstruction Network
 1 Introduction
-The rise of the internet and digital technologies has significantly accelerated information dissemination while also amplifying the spread of misinformation, enabling new forms of deceptive content such as DeepFakes [1], multimodal misinformation [2], and LLM-generated misinformation [3].
-Given the scale and rapid spread of misinformation, researchers have been developing automated fact-checking methods to support human fact-checkers in identifying false information more efficiently [4].
-In this study, we focus on multimodal misinformation detection (MMD), specifically, the detection of misleading image-caption pairs, where texts and images jointly contribute to the spread of false or deceptive information [5].
-Recent studies on MMD primarily focus on developing large-scale datasets and detection methods.
-Aside from a few small-scale annotated datasets [6, 7], existing datasets primarily consist of weakly annotated [8, 9, 10] or algorithmically generated datasets; created either by pairing images with out-of-context (OOC) captions from other images [11, 12, 13, 14] or by manipulating named entities to introduce inconsistencies, resulting in miscaptioned (MC) images [15, 16, 17].
-On the modeling side, detection models have leveraged large pre-trained backbone encoders [14] or fine-tuned through self-supervised learning [18], improved modality fusion with attention-based techniques [19, 20, 21], incorporated external evidence [22, 23], and utilized Large Vision-Language Models (LVLMs) for both detection [24, 25] and explanation generation [26].
-Nevertheless, named-entity manipulations often produce simplistic misinformation, lacking basic factual or logical consistency, while the potential of LVLMs to generate more diverse and robust synthetic training data remains unexplored.
-Moreover, despite the success of reconstruction networks in other domains, their potential remains largely unexplored for MMD where is particularly relevant, as the task closely aligns with human fact-checking practices, where reconstructing the true origin, context, or meaning of an image is a common strategy for debunking misinformation [27].
-Figure 1: High-level overview of the proposed framework: an LVLM modifies a truthful image-caption pair, while LAMAR’s reconstruction network re-creates the original caption embedding produced by CLIP.
-This embedding is fused with other modalities using a mechanism (e.g., Gate) and fed into the detection network to produce the final verdict.
-The reconstruction network is trained to minimize the error between original and reconstructed embeddings, while the detection network is optimized for classification.
-To this end, we propose a framework that leverages an LVLM to manipulate the original captions of images sourced from a dataset of truthful image-caption pairs, generating false captions that misrepresent aspects of the images.
-Subsequently, the proposed Latent Multimodal Reconstruction (LAMAR) network is tasked with reconstructing the embeddings of the original, truthful captions.
+The rise of the Internet and digital technologies has significantly accelerated the spread of information but also the proliferation of misinformation, including new forms of deceptive content such as DeepFakes [32], multimodal misinformation [2], and LLM-generated misinformation [8].
+Given the scale and speed of misinformation dissemination, researchers are developing automated fact-checking tools to assist human fact-checkers in identifying deceptive content more efficiently [11].
+In this study, we focus on multimodal misinformation detection (MMD), specifically, the detection of misleading image-caption pairs, where both modalities jointly contribute to the spread of false or misleading information [3].
+Due to the scarcity of large-scale, expert-annotated MMD datasets, researchers have increasingly turned to synthetic training data.
+While fact-checking archives provide high-fidelity labels, their limited volume of multimodal pairs, lack of negative samples, and licensing constraints render them useful for evaluation but insufficient for training.
+Consequently, aside from a few small-scale annotated resources [44], recent studies have had to rely on either weakly annotated [6, 25, 26, 15] or algorithmically generated training datasets; created either by pairing images with out-of-context (OOC) captions from other images [13, 4, 5, 22] or by manipulating named entities to introduce inconsistencies, resulting in miscaptioned (MC) images [33, 24, 30].
+However, named-entity manipulation often produces shallow or implausible misinformation, lacking the nuance seen in real-world examples.
+At the same time, the potential of Vision Language Models (VLMs) to generate richer, more diverse synthetic training data remains underexplored.
+On the modeling side, recent works have explored large pre-trained encoders [22], self-supervised fine-tuning [23], attention-based fusion mechanisms [18, 40, 27], external evidence [1, 41], and Vision-Language Models (VLMs) for both detection [31, 35] and explanation generation [42].
+However, while reconstruction networks have shown promise in other domains, their use in MMD is still limited – despite their alignment with human fact-checking strategies, where reconstructing the original meaning, context, or source of an image is key to detecting misinformation [7].
+Figure 1: High-level overview of the proposed workflow.
+A VLM generates a false caption from the original, truthful image-caption pair.
+The Reconstructor (Transformer encoder) then takes both the image and false caption embeddings as inputs to identify and rectify inaccuracies in the text, recreating the original truthful caption embedding.
+This reconstructed representation is fused with the other modalities via a specialized mechanism (e.g., Gating or Attention) and passed to the detector to produce the final verdict.
+The network is trained to simultaneously minimize the error between original and reconstructed embeddings and optimized for classification accuracy.
+To this end, we propose “MisCaption This!”, a framework that leverages VLMs to manipulate the captions of images drawn from truthful image-caption datasets, generating synthetic false captions that misrepresent aspects of the image.
+The resulting datasets are used to train the proposed Latent Multimodal Reconstruction (LAMAR) network, which learns to reconstruct the embedding of the original, truthful caption.
 As shown in Fig.
-1, a truthful image-caption pair is manipulated by the LVLM, while LAMAR is tasked with reconstructing the original caption embedding using the image and the generated caption embeddings.
-Thereafter, the reconstructed embedding is integrated into the final detection network alongside the fused modalities.
-Our rationale is that LVLMs can generate more diverse and realistic training data, enhancing the generalizability of detection models, while the reconstruction process serves as an auxiliary signal, leveraging learned representations to refine embeddings by capturing consistency and inconsistency patterns in image-caption pairs—even in the absence of external evidence.
-More specifically, we leverage an LVLM alongside “adversarial prompt selection” to limit the number of generative prompts by evaluating them against the LVLM’s zero-shot detection capabilities; filtering out prompts that produce easily detectable misinformation or overly generic image descriptions.
-This process results in the selection of four generative prompts, each used to generate a distinct version of our “MisCaption This!” dataset.
-For LAMAR, we employ a Transformer encoder with element-wise vector operations for modality fusion and explore two training strategies, end-to-end training and large-scale pre-training, as well as four methods for integrating the reconstructed embeddings into the detection network: direct integration, masking, gating, and self-attention.
-Through extensive experiments with various training datasets, we demonstrate that models trained on “MisCaption This!” achieve superior out-of-distribution generalization on the VERITE evaluation benchmark [28], outperforming those trained on datasets relying on entity manipulation or cross-modal misalignment—with LAMAR performing 7.8% and 10.4% better, respectively.
-Moreover, LAMAR consistently outperforms prior state-of-the-art (SotA) models across training settings and datasets, achieving improvements of 4.3% on ‘True vs.
-MC’, 3.0% on ‘True vs.
-OOC’, and 4.8% on the multiclass task when trained on “MisCaption This!”.
+1, LAMAR takes the image and manipulated caption embeddings as input and is tasked with recovering the original caption embedding.
+This reconstructed embedding is then integrated into the final detection network alongside the fused multimodal representation.
+Our rationale is that (1) VLMs can generate more diverse and realistic synthetic training data, and (2) semantic divergence between reconstructed and input embeddings provides a discriminative signal that exposes cross-modal and intra-modal inconsistencies, significantly improving model generalization.
+“MisCaption This!” uses open-source VLMs (LLaVA, Llama, Molmo) and apply adversarial prompt selection, where generative prompts are evaluated against the VLM’s zero-shot detection capabilities.
+Prompts that produce easily detectable or overly generic captions are discarded.
+LAMAR employs a Transformer encoder with element-wise modality fusion, and we explore both end-to-end training and large-scale pre-training.
+We also compare four integration strategies for the reconstructed embedding: direct integration, masking, gating, and self-attention.
+Our main contributions can be summarized as follows:
+  • We introduce “MisCaption This!”, a framework for creating synthetic training datasets of miscaptioned images using open-source VLMs.
+  • We propose the Latent Multimodal Reconstruction (LAMAR) architecture, which learns to reconstruct the truthful caption embedding from its input.
+  • We show that models trained on VLM-generated data generalize better to real-world misinformation, outperforming named-entity swapping and cross-modal misalignment datasets by 7.8% and 10.4% on the VERITE benchmark [28], respectively.
+  • LAMAR achieves new state-of-the-art (SotA) results on NewsCLIPpings and across all VERITE tasks, with improvements of 4.3% on “True vs.
+MC”, 3.0% on “True vs.
+OOC”, and 5.6% on multiclass classification.
+  • LAMAR demonstrates stronger temporal generalization and robustness on emerging misinformation compared to zero-shot VLMs, maintaining high performance (70.0%) on the new VERITE 24/25 benchmark – consisting of events after the VLM knowledge-cutoff – while VLM performance degrades significantly (8–12% reduction).
 2 Related Work
-In recent years, automated fact-checking has attracted growing research interest [4], encompassing a range of challenging tasks such as claim detection [29], evidence filtering and retrieval [30, 31], fake news detection [32, 33, 34], retrieval of fact-checked articles [35, 36], DeepFake detection [1], and multimodal forms of misinformation [2].
-In this paper, we focus on MMD, which is garnering increasing attention from researchers, leading to the ongoing development of datasets and methodologies for detecting false information and inconsistencies between images and their accompanying textual captions.
 2.1 MMD Datasets
 Training machine learning models for MMD requires suitable datasets, with current research focusing on annotated, weakly annotated, and synthetically generated data.
-Early MMD datasets, such as the ‘Twitter’ [7] and ‘Weibo’ [37] datasets, are relatively small and cover only a limited number of events-17 and 73, respectively-raising concerns about model generalization.
+Early MMD datasets, such as the Twitter [6] and Weibo [14] datasets, are relatively small and cover only a limited number of events-17 and 73, respectively-raising concerns about model generalization.
 To address this, larger weakly annotated datasets have emerged, including MuMiN [26], with rich social context but few images, and NewsBag [15], which includes satirical content, and Fakeddit [25], with over a million instances collected from Reddit.
-However, studies show that models trained on ‘Twitter’ and Fakeddit often exhibit unimodal biases, undermining their effectiveness in real-world multimodal misinformation detection [28].
+However, studies show that models trained on Twitter and Fakeddit often exhibit unimodal biases, undermining their effectiveness in real-world multimodal misinformation detection [28].
 Researchers have also explored synthetic data generation.
@@
 NES-based datasets generate misinformation by replacing named entities in captions with alternatives retrieved from similar or contextually relevant texts using cluster-based retrieval (MEIR [33]), rule-based substitutions (TamperedNews [24]), and CLIP-based retrieval (CLIP-NESt [30]).
-Despite advancements in LVLMs, only MMFakeBench leverages LVLM-generated rumors and AI-altered images and serves as a small evaluation benchmark [38].
-Instead, we introduce a large LVLM-generated training dataset.
+Despite recent progress in VLMs and their use for synthetic dataset generation [19], their application to MMD remains limited.
+To date, only MMFakeBench [21] uses VLM-generated rumors and AI-manipulated images, and serves solely as a small-scale evaluation benchmark.
+In contrast, we introduce a large-scale VLM-generated training dataset for MMD.
 However, models trained and evaluated on synthetic data may struggle with real-world generalizability, as they may learn to detect patterns specific to artificially generated inconsistencies rather than the more complex and diverse manipulations found in real-world misinformation.
@@
 Research on MMD has centered on developing models that encode textual and visual modalities, fuse their representations, and assess their consistency and factual accuracy.
-Early methods, such as SpotFake [39], used VGG-19 and BERT, while recent approaches use pre-trained multimodal encoders such as CLIP [14], or fine-tune it through self-supervised learning [18].
+Early methods, such as SpotFake [34], used VGG-19 and BERT, while recent approaches use pre-trained modality-specific dual-encoder models such as CLIP [22], or fine-tune it through self-supervised learning [23].
 Some models incorporate multi-task learning, such as EANN with an event discriminator [36] or MVAE, which uses an autoencoder to reconstruct the input text and visual features [17]; but does not modify the input text or reconstruct truthful captions from false ones.
@@
 Recent work integrates external web evidence, with methods assessing internal and external consistency (CCN [1], SNIFFER [31]) or evaluating stance and relevance of external evidence (SEN [41], RED-DOT [27]).
-While external evidence is shown to improve performance, concerns remain about ‘leaked evidence’ from fact-checking articles [42, 31] and dataset artifacts that models may exploit instead of assessing factuality [43].
+While external evidence is shown to improve performance, concerns remain about ‘leaked evidence’ from fact-checking articles [10, 9] and datasets-specific artifacts that models may exploit instead of assessing factuality [29].
 For these reasons, we do not consider evidence-based approaches in this study.
@@
 Reconstruction networks remain largely unexplored in MMD, with MVAE [17] being a notable exception.
-However, MVAE uses an autoencoder to reconstruct the input text and image embeddings from a joint latent space but does not modify the text or generate truthful image descriptions from false ones; its objective is to extract more informative features for MMD within the latent space.
+However, MVAE employs a variational autoencoder to reconstruct the input, recovering the input words (via Bi-LSTMs) and VGG-19 image embeddings from a shared latent space.
+Crucially, this reconstruction serves only to enhance feature extraction and does not involve modifying or correcting the input content, e.g., generating truthful descriptions, or their corresponding latent representations, from false ones.
 3 Problem Formulation
-Given a set $\mathcal{D}^{t}={(I^{t}_{i},C^{t}_{i})}_{i=1}^{N}$ of $N$ image-caption pairs, where $I^{t}_{i}$ and $C^{t}_{i}$ represent an image and its matching, truthful caption, we define ‘Out-Of-Context’ (OOC) as any pair that combines an image $I^{t}_{i}$ with a caption $C^{x}_{i}$ taken from another image within $\mathcal{D}^{t}$; where x denotes the context.
-Similarly, we define ‘Mis-Captioned’ (MC) as any image whose accompanying caption $C^{f}_{i}$ has been manipulated to misrepresent the original, content, and/or meaning of the image; where “f” denotes falsehood.
-We define a “Manipulator” as any method used to generate OOC pairs (e.g., CLIP-based retrieval) or MC pairs (e.g., named entities manipulation) from the original set $\mathcal{D}^{t}$ set, resulting in the sets $\mathcal{D}^{x}$ and $\mathcal{D}^{f}$, respectively.
-We define MMD as a classification task with the objective to learn a mapping function $\mathsf{M}^{d}:\mathcal{D}\rightarrow\hat{y}$ where $\hat{y}$ represents the prediction upon target class $y$ in either of three scenarios:
-1) Binary: ‘True vs. MC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{f}]$ with $K=N*2$ total pairs and $y\in\{0,1\}$
-2) Binary: ‘True vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x}]$ with $K=N*2$ total pairs and $y\in\{0,2\}$,
-3) Multi-class: ‘True vs. MC vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x},\mathcal{D}^{f}]$ with $K=N*3$ total pairs and $y\in\{0,1,2\}$.
-Given image and text encoders $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$ producing image embeddings $\mathbf{I}$ and text embeddings $\mathbf{C}$ for image-caption pairs $(I,C)$ under examination, we define latent reconstruction as the task of learning a mapping function $\mathsf{M}^{r}:(\mathbf{I},\mathbf{C})\rightarrow\mathbf{\hat{C}^{t}}$, where $\mathbf{\hat{C}^{t}}$ represents the predicted embedding of the original, truthful caption $\mathbf{C^{t}}$.
-The model $\mathsf{M}^{r}$ is trained to minimize the reconstruction error between the true and predicted caption embeddings, such that:
+Given a set $\mathcal{D}^{t}={(I^{t}_{i},C^{t}_{i})}_{i=1}^{N}$ of $N$ image-caption pairs, where each $I^{t}_{i}$ is an image and $C^{t}_{i}$ its matching, truthful caption, we define:
+Out-of-Context (OOC) pairs as $(I^{t}_{i},C^{x}_{i})$, where $C^{x}_{i}$ is a caption taken from a different sample in $\mathcal{D}^{t}$;
+Mis-Captioned (MC) pairs as $(I^{t}_{i},C^{f}_{i})$, where $C^{f}_{i}$ is a manipulated version of the original caption that misrepresents the content or meaning of $I^{t}_{i}$.
+A “manipulator” refers to any method used to create OOC (e.g., CLIP-based retrieval) or MC (e.g., entity swapping) samples, producing datasets $\mathcal{D}^{x}$ and $\mathcal{D}^{f}$ from $\mathcal{D}^{t}$.
+We define MMD as a classification task to learn a mapping $\mathsf{M}^{d}:\mathcal{D}\rightarrow\hat{y}$, where $\hat{y}$ is the predicted class label under one of three settings:
+(1) Binary: ‘True vs. MC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{f}]$ with K=N2K=N*2 total pairs and $y\in\{0,1\}$
+(2) Binary: ‘True vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x}]$ with K=N2K=N*2 total pairs and $y\in\{0,2\}$,
+(3) Multi-class: ‘True vs. MC vs. OOC’, where $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{x},\mathcal{D}^{f}]$ with K=N3K=N*3 total pairs and $y\in\{0,1,2\}$.
+Let $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$ denote the image and text encoders, producing embeddings $\mathbf{I}$ and $\mathbf{C}$ for input pair $(I,C)$, respectively.
+We define latent reconstruction as learning a function $\mathsf{M}^{r}:(\mathbf{I},\mathbf{C})\rightarrow\hat{\mathbf{C}}^{t}$, where $\hat{\mathbf{C}}^{t}$ approximates the embedding of the original truthful caption $\mathbf{C}^{t}$.
+The model is trained to minimize reconstruction loss:
 $\mathcal{L}_{r}(C^{t},\hat{C}^{t})=\frac{1}{K}\sum_{i=1}^{K}(\mathbf{C}^{t}_{i}-\hat{\mathbf{C}}^{t}_{i})^{2}$ (1)
-where $\mathcal{L}_{r}$ is the loss function (i.e., the Mean Squared Error (MSE)) that quantifies the discrepancy between the embeddings of the true and reconstructed caption embeddings.
-Figure 2: Examples of truthful and generated captions taken from the “MisCaption This!” ($\mathcal{D}_{3}$) dataset, alongside false captions created via named entity swaps.
-4 Construction of “MisCaption This!”
-In this study, we explore the creation of a synthetic training dataset of miscaptioned images ($\mathcal{D}^{f}$) by manipulating the image captions of a truthful dataset ($\mathcal{D}^{t}$) using an LVLM as the “Manipulator”.
-Our rationale is that LVLMs, with their advanced multimodal understanding and generation capabilities, can produce more realistic false captions for images compared to methods relying on manipulating named entities.
-In turn, we hypothesize that the generated data $\mathcal{D}^{f}$ can be leveraged to train more robust detection models $\mathsf{M}^{d}$, thus enhancing their ability to generalize to real-world misinformation.
+where $\mathcal{L}_{r}$ is the Mean Squared Error (MSE) between the true and reconstructed caption embeddings.
+Unless explicitly denoted as $C^{t}$ (truthful), $C^{f}$ (falsified), or $\mathbf{\hat{C}}^{t}$ (reconstructed), we use $C$ and $\mathbf{C}$ to refer to the input caption and its embedding–regardless of truthfulness.
+4 “MisCaption This!”
+In this study, we explore the creation of synthetic training datasets of miscaptioned images ($\mathcal{D}^{f}$) by manipulating the image captions of a truthful dataset ($\mathcal{D}^{t}$) using a VLM as the “Manipulator”.
+Our rationale is that VLMs, with their advanced multimodal understanding and generation capabilities, can produce more realistic false captions for images compared to methods relying on manipulating named entities.
+In turn, we hypothesize that the generated data $\mathcal{D}^{f}$ can be leveraged to train more robust detection models $\mathsf{M}^{d}$, thus enhancing generalization to real-world misinformation.
 4.1 Generative Model
-To generate synthetic data, we employ LLaVa-1.6 (Large Language and Vision Assistant) [48] from Hugging Face [^1] https://huggingface.co/llava-hf/llava-v1.6-mistral-7b-hf leveraging Mistral-7B-Instruct-v0.2.
-We also explored LLaVa-1.6 (Vicuna-13B), Janus Pro 7B (DeepSeek), and MiniGPT-v2 (Llama-2-7B) as alternatives, however, we were unable to get them to consistently generate realistic false captions, as they often defaulted to generic image captioning, re-phrasing of the original caption, or overly simplistic misinformation.
+To generate synthetic training data, we employ open-source VLMs: LLaVA-1.6-Mistral [^1] https://huggingface.co/LLaVA-hf/LLaVA-v1.6-mistral-7b-hf, Molmo-7B-D [^2] https://huggingface.co/allenai/Molmo-7B-D-0924, and Llama-3.2-11B-Vision [^3] https://huggingface.co/meta-llama/Llama-3.2-11B-Vision.
+We also explored LLaVA-1.6-Vicuna-13B, Janus Pro 7B (DeepSeek), and MiniGPT-v2-Llama-2-7B as alternatives, however, we were unable to get them to consistently generate realistic false captions, as they often defaulted to generic image captioning, re-phrasing of the original caption, or overly simplistic misinformation.
 In contrast, GPT-4o Mini demonstrated robust safeguards, often refraining from generating misinformation altogether.
-Therefore, this part of our work serves as an initial exploration of LVLM-generated synthetic data, with future research needed to assess the impact of other LVLMs and prompting methods.
 4.2 Adversarial Prompt Selection
-Generating realistic misinformation with an LVLM requires a carefully selected generative prompt $p^{gen}$.
-However, determining which $p^{gen}$ will produce a dataset $\mathcal{D}^{f}$ that leads to better-performing detection models $\mathsf{M}^{d}$ is not straightforward and would require significant resources for an extensive empirical examination.
-To address this, we propose Adversarial Prompt Selection, which refines the set of generative prompts by assessing their generated captions $C^{f}$ based on their ability to evade detection by a zero-shot LVLM, ensuring that the generated data presents a meaningful challenge for detection models, rather than being trivially misleading or overly simplistic.
-First, we evaluate the LVLM’s ability to perform zero-shot detection of miscaptioned images.
-We experiment with multiple detection prompts for LLaVa, We experiment with multiple detection prompts for LLaVa, varying their formulation (e.g., word synonyms, rephrasing) and compared their zero-shot performance on a subset of 100 real-world “True vs.
-MC” samples from VERITE.
-Based on this process, the detection prompt ($p^{dt}$): “The image is captioned as: [CAPTION].
-Is the caption truthful or does it contain falsehoods?” was selected for the next stage.
-Next, we evaluate a range of generative prompts against zero-shot LLaVa with $p^{dt}$, experimenting with various formulations similar to those used for the detection prompts.
-For each generative prompt ($\forall j\in J$), LLaVa generates 1,000 captions from a randomly selected subset of 1,000 truthful image-caption pairs from the NewsCLIPpings training set.
-Then, the LVLM is tasked with classifying both the truthful and generated captions, and we compute the average detection accuracy for each prompt.
-This process can be formulated as:
-$C^{f}_{j}=\text{LLaVa}(I,C^{t}\mid p^{gen}_{j}),\quad\forall j\in[1,J]$ (2)
-$\hat{y}_{j}=\text{LLaVa}(I,C_{j}\mid p^{dt}),\quad C_{j}\in\{C^{t}_{j},C^{f}_{j}\}$ (3)
-We reason that generated captions produced by simplistic prompts, which are easily detected, indicate the generation of unrealistic misinformation and are unlikely to contribute to training effective $\mathsf{M}^{d}$ models.
-For instance, generated captions from prompts such as ‘Create a false caption for the image to propagate misinformation’ and ‘Create fake news based on the image’ were detected with 83.5% and 81.0% accuracy, respectively.
-Conversely, generative prompts that produce undetectable captions, may fail to produce actual misinformation, instead leaving the original captions unchanged or rephrased.
-For example, the prompts ‘Provide a caption for the image’ and ‘Maintain the original caption of the image’ resulted in low detection accuracy (42.5% and 50.0%) while having low (7%) or perfect (100%) word overlap with the original captions, indicating a generic or truthful image description rather than misinformation.
-Through this process, and after manual inspection, we select four generative prompts, $p^{gen}_{1}$, $p^{gen}_{2}$, $p^{gen}_{3}$, and $p^{gen}_{4}$, with varying adversarial zero-shot detection performance: 57%, 62%, 71%, and 75%.
-These prompts are then used to generate four distinct sets, $\mathcal{D}^{f}_{1}$, $\mathcal{D}^{f}_{2}$, $\mathcal{D}^{f}_{3}$, and $\mathcal{D}^{f}_{4}$, each corresponding to a different set of false captions related to the images of $\mathcal{D}^{t}$.
-All four prompts are multimodal, as they take into account both the image and its original caption to generate false captions.
-For ethical considerations, we refrain from disclosing the specific prompts here, as sharing them could contribute to the automated generation of misinformation; however, they will be available upon request under a research-only license.
+Generating realistic misinformation with an VLM requires carefully selected generative prompts ($p^{gen}$).
+However, identifying an “optimal” prompt for a task is non-trivial and costly to evaluate exhaustively.
+To mitigate this, we leverage Adversarial Prompt Selection, which filters prompts based on how well their generated captions ($C^{f}$) evade detection by a zero-shot VLM, ensuring the generated data presents a meaningful challenge.
+First, we assess the VLM’s (LLaVA) zero-shot capability as a ‘Detector’ of miscaptioned images.
+We conduct a grid search over a candidate pool of prompts varying in wording and granularity.
+These were evaluated on a balanced calibration set ($N=100$) from the VERITE (“True vs.
+MC”) benchmark.
+Through this process, we identified the optimal prompt as: “The image is captioned as: [CAPTION].
+Is the caption truthful or does it contain falsehoods?”, which achieved $\approx 70\%$ accuracy.
+This prompt was selected for subsequently experiments as the adversary detection prompt ($p^{dt}$) against which we evaluate candidate generative prompts ($p^{gen}$).
+Our analysis revealed that the specific phrasing “contain falsehoods” consistently outperformed synonyms such as “manipulated”, “fake”, or “falsified”.
+Furthermore, explicitly defining terms, such as stating that false captions “may falsely describe origin, context, and/or meaning of the image” or “contain inconsistencies between named entities (people, dates, locations etc.)” hindered performance, yielding only $\approx 62\%$ accuracy.
+Figure 2: Adversarial Prompt Selection: An VLM ‘Manipulator’ generates a false caption using the generative prompt $p_{0}^{gen}$.
+The VLM ‘Detector’ is then evaluated on the zero-shot classification of both the truthful and generated captions.
+Intermittent lines indicate the prediction on a single sample, while the overall accuracy is calculated across a balanced set of 2,000 samples.
+In this specific case, the high overall detection accuracy indicates the generated misinformation is too simplistic; thus $p_{0}^{gen}$ is not selected for the creation of a training dataset.
+Next, we assess a range of generative prompts by using the VLM as ‘Manipulator’ to produce 1,000 falsified captions ($C^{f}_{j}$) for 1,000 randomly sampled truthful examples from NewsCLIPpings, as expressed in Eq.
+2.
+Generated captions ($C^{f}_{j}$) are assigned the ground truth label ‘False’.
+To evaluate the quality of the generated misinformation, we then subject both the generated captions ($C^{f}_{j}$) and the original, truthful captions ($C^{t}$ - labeled ‘True’) to zero-shot classification by the VLM-Detector using the selected detection prompt, as expressed in Eq.
+3.
+$C^{f}_{j}=\text{VLM-Manipulator}(I,C^{t}\mid p^{gen}_{j}),\quad\forall j\in[1,J]$ (2)
+$\hat{y}_{j}=\text{VLM-Detector}(I,C_{j}\mid p^{dt}),\quad C_{j}\in\{C^{t}_{j},C^{f}_{j}\}$ (3)
+Prompts that result in high detection accuracy often generate simplistic or unrealistic misinformation.
+For example, as seen in Fig.
+2, the prompt “Create a false caption for the image to propagate misinformation” produced: “UFO sighting: Alien spacecraft descends on Iranian airbase, preparing for an intergalactic invasion” for this specific image, and the VLM-Detector yielded high detection accuracy (83.5%).
+Conversely, prompts yielding very low detection accuracy could mean that the generated captions are truthful, generic or re-phrased captions, failing to introduce actual falsehoods.
+For example, the generic prompt “Provide a caption for the image” produced “A lone fighter jet soaring through the clear blue sky, leaving a trail of contrails behind” for the same image, with low overall detection accuracy (42.5%).
+Therefore, we do not select prompts that produce either easily detectable false captions (due to being unrealistic) or those that fail to introduce falsehoods (due to either being truthful or generic descriptions).
+Through this process and manual inspection, we select four generative prompts ($p^{gen}_{1}$$p^{gen}_{4}$) that span a range of adversarial difficulty (57%, 62%, 71%, 75% detection accuracy).
+These are used to generate four datasets ($\mathcal{D}^{f}_{1}$$\mathcal{D}^{f}_{4}$) of falsified captions conditioned on both the image and its original caption.
+After identifying the most effective generative prompt ($p^{gen}_{3}$) for LLaVA, we use it for Molmo and Llama as well to keep the computational load of our experiments to reasonable levels while testing cross-model generalizability.
+For ethical reasons, we withhold exact prompts but make them available upon request under a research-only license.
+Figure 3: Examples of truthful and generated captions, alongside false captions created via named entity swaps.
 4.3 Generated Data Filtering
-Through manual examination, we observe that generative prompts produce more coherent and plausible false captions, often better than NES-based methods in terms of logical coherence and knowledge-based consistency.
-For instance, as shown in Fig.2, manipulating named entities led to noticeable inconsistencies, such as the incorrect claims of the “Indian parliament” being in “Bangladesh” and “Skeldergate” being in “San Francisco”.
-These examples underscore the limitations of NES-based methods in preserving basic logical and factual consistency.
-In contrast, the LVLM demonstrated the ability to generate more creative and plausible misinformation.
-For instance, it misrepresented a demonstration by reframing its purpose, shifting it from anti-austerity to pro-austerity, and further reinforced the false narrative with claims of “successful implementation” and “widespread approval”.
-Similarly, it distorted the aftermath of the evacuation following heavy flooding, suggesting that people were “enjoying a leisurely swim in the flooded streets” and “turning a disaster into a fun-filled day”.
-However, we also observed that LLaVa can occasionally “ramble”, often rephrasing the original caption while adding redundant or superficial details.
-For instance, given the original caption “The recent economic boom has enabled new projects such as the Union Trade Centre shopping centre in the heart of Kigali”, LLaVa generated: “The Union Trade Centre shopping centre in Kigali is a prime example of the city’s thriving economy with numerous cars parked outside indicating a bustling shopping scene”; thereby rephrasing the original caption and incorporating minor descriptive elements that do not amount to misinformation.
-To address this issue, we apply a post-processing filter that removes samples where the length of the generated caption exceeds a specified character threshold $l\in{\{0,5,10,15,25,50,None\}}$ relative to the original, truthful caption.
-To maintain a balanced dataset, we remove both the generated pairs ($I^{t}_{i},C^{f}_{i}$) and their corresponding truthful pairs ($I^{t}_{i},C^{t}_{i}$) that exceed this threshold.
-This filtering process retains 4.5%, 19.1%, 27.8%, 34.9%, 47.9%, 74.0%, and 100% of the dataset for $l\in{\{0,5,10,15,25,50,None\}}$, respectively.
-We empirically evaluate the impact of this filtering on model performance.
+Qualitative analysis (such as in Fig.
+3) showed that VLMs can produce more coherent and plausible false captions than NES-based methods, especially in terms of logical and factual consistency.
+For instance, NES manipulations yielded implausible claims like the ‘Indian parliament’ being in ‘Bangladesh’ or ‘Skeldergate’ in ‘San Francisco’, highlighting the limits of surface-level entity substitution.
+In contrast, VLMs showed the ability to generate more nuanced and believable misinformation.
+It might, for instance, reframe an anti-austerity protest as pro-austerity, reinforcing the false narrative with phrases like “successful implementation” and “widespread approval”, or distort the context of flood evacuations by suggesting that people were “enjoying a leisurely swim” and “turning a disaster into a fun-filled day”.
+Nevertheless, LLaVA occasionally “rambles”, rephrasing the original caption with superficial elaborations rather than introducing actual misinformation.
+For instance, from the original caption: “The recent economic boom has enabled new projects such as the Union Trade Centre shopping centre in the heart of Kigali”, LLaVA generated: “The Union Trade Centre shopping centre in Kigali is a prime example of the city’s thriving economy with numerous cars parked outside indicating a bustling shopping scene”, which merely rephrases the original with extraneous details.
+Moreover, if generation often results in longer captions, this could introduce superficial patterns (“shortcuts”), artifacts, that detection models may learn to exploit, instead of learning deeper features that are associated with content factuality and credibility.
+To address this, we apply a post-processing filter that excludes samples where the generated caption exceeds the original by a relative character-length threshold $l\in\{0,5,10,15,25,50,\text{None}\}$, retaining 4.5%, 19.1%, 27.8%, 34.9%, 47.9%, 74.0%, and 100% of the LLaVA-$\mathcal{D}_{3}$ dataset, respectively.
+We empirically evaluate how filtering affects model performance.
+To maintain class balance, we removed both the generated pair $(I^{t}_{i},C^{f}_{i})$ and its corresponding truthful pair $(I^{t}_{i},C^{t}_{i})$ when this threshold is exceeded.
 4.4 Dataset Source and Statistics
-We use the NewsCLIPpings ‘Merged/Balanced’ version [14], which has been shown to be effective for OOC detection [22, 24, 21, 43], as the source dataset $\mathcal{D}^{t}$ from which to generate $\mathcal{D}^{f}$, utilizing only its truthful data: 35,536 pairs for training, 3,512 for validation, and 3,512 for testing.
+We use the NewsCLIPpings ‘Merged/Balanced’ dataset [22], which has been shown to be effective for OOC detection [1, 31, 27, 29], as the source dataset $\mathcal{D}^{t}$ from which to generate $\mathcal{D}^{f}$, utilizing only its truthful data: 35,536 pairs for training, 3,512 for validation, and 3,512 for testing.
 After generating $\mathcal{D}^{f}_{1}$, $\mathcal{D}^{f}_{2}$, $\mathcal{D}^{f}_{3}$, and $\mathcal{D}^{f}_{4}$ from $\mathcal{D}^{t}$ using LLaVA and the corresponding generative prompts, we merge $\mathcal{D}=[\mathcal{D}^{t},\mathcal{D}^{f}]$ to address the “True vs.
@@
 This results in a total of 106,605, 10,536, and 10,896 samples for training, validation, and testing, respectively, ensuring a balanced distribution across the three classes.
+We use the Molmo and Llama generated datasets in the same manner.
 5 Latent Multimodal Reconstruction (LAMAR)
-Our objective is to develop a reconstruction network $\mathsf{M}^{r}$ that utilizes the embeddings $\mathbf{I},\mathbf{C}$ of image-caption pairs to reconstruct the embedding of the original, truthful caption associated with the image.
-The reconstructed embeddings $\mathbf{\hat{C}^{t}}$ are then integrated within the final detection model, along with the fused modalities, to aid the detection process.
-Our rationale is that if the reconstructed embedding $\mathbf{\hat{C}^{t}}$ closely match the input embeddings $\mathbf{C}$, then, the pair $\mathbf{I},\mathbf{C}$ is likely to be truthful.
-On the other hand, significant discrepancies between $\mathbf{C}$ and $\mathbf{\hat{C}^{t}}$ should suggest potential manipulation of $C$.
-We hypothesize that this process will provide valuable signals for the detection model.
-Since reconstruction networks remain under-explored in the context of MMD, we investigate various alternative strategies, broadly categorized into: (1) end-to-end training, and (2) large-scale pre-training.
-Additionally, we explore attention, gating, and masking mechanisms to refine the integration of the reconstructed embedding into the Detection network.
-As illustrated in Figure 3, our pipeline consists of a backbone encoder, followed by modality fusion, a transformer-based reconstruction module, integration of the reconstructed embedding, and a final classification layer.
-We provide a detailed discussion of these components in the following sections.
-Figure 3: End-to-end training of the proposed LAMAR architecture, which utilizes a CLIP ViT L/14 backbone encoder and a Transformer encoder for the reconstruction network with element-wise vector operations for enhanced modality fusion.
-The reconstructed embedding is then integrated into the detection network through various mechanisms (e.g., gate, mask, attention, or no mechanism) which predicts the final verdict.
-The reconstruction network is optimized using MSE loss to minimize the difference between reconstructed and ground truth caption embeddings, while the detection network is optimized with cross-entropy (CE) loss.
+We propose a reconstruction module $\mathsf{M}^{r}$ that uses image-caption embeddings $(\mathbf{I},\mathbf{C})$ to reconstruct the embedding of the original truthful caption $\hat{\mathbf{C}}^{t}$.
+The reconstructed embedding serves as an additional signal for the detection model: alignment between $\hat{\mathbf{C}}^{t}$ and the input $\mathbf{C}$ indicates a likely truthful pair, while significant divergence may suggest manipulation.
+We explore two training strategies—end-to-end and large-scale pre-training—and investigate mechanisms like attention, gating, and masking for integration.
+As shown in Fig.
+4, including a backbone encoder, modality fusion, reconstruction, and detection modules.
+Figure 4: End-to-end training of the proposed LAMAR architecture.
+The input caption $C$ may be either truthful or falsified.
+A CLIP ViT-L/14 encoder and a Transformer-based reconstruction module, enhanced with element-wise vector operations for modality fusion, outputs the reconstructed embedding $\mathbf{\hat{C}^{t}}$.
+This is integrated into the detection network via some mechanisms (e.g., gate, mask, attention), which outputs the final verdict.
+The reconstruction module is trained with MSE loss against the ground-truth $\mathbf{C}^{t}$, while the detection module is trained with cross-entropy (CE).
 5.1 Backbone Encoder
-We use CLIP ViT L/14 from OpenCLIP  [^2] https://github.com/mlfoundations/open_clip as the backbone multimodal encoder, $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$, to produce image embeddings $\mathbf{I}\in{R}^{768\times 1}$ and text embeddings $\mathbf{C}\in{R}^{768\times 1}$, which are pre-aligned within a shared embedding space.
+We use CLIP ViT L/14 from OpenCLIP  [^4] https://github.com/mlfoundations/open_clip as the backbone multimodal encoder, $\mathsf{E}_{I}$ and $\mathsf{E}_{C}$, to produce image embeddings $\mathbf{I}\in\mathbb{R}^{768\times 1}$ and text embeddings $\mathbf{C}\in\mathbb{R}^{768\times 1}$, which are pre-aligned within a shared embedding space.
 5.2 Modality Fusion
@@
 with $\mathbf{F}\in{R}^{5\times 768}$.
+Concatenation $\mathbf{I};\mathbf{C}]$ provides the original, isolated features, while the element-wise operations $\mathbf{I}+\mathbf{C}$, $\mathbf{I}-\mathbf{C}$, and $\mathbf{I}*\mathbf{C}$ are included to highlight complementarity, differences, and mutual agreement between the visual and textual embeddings, respectively.
 5.3 Detection Network
@@
 Bias terms $b$ are included in the model but omitted here for brevity.
-Here, the operation ‘Flatten’ refers to converting the concatenated vector $[\mathbf{F};\mathbf{\hat{C}^{t}}]$ into a one-dimensional vector before passing it through the hidden layers.
+Here, the operation ‘Flatten’ refers to converting the concatenated vector $\mathbf{F};\mathbf{\hat{C}^{t}}]$ into a one-dimensional vector before passing it through the hidden layers.
 The Detection Network is optimized using binary cross-entropy loss for binary classification and categorical cross-entropy for multi-class classification; denoted as $\mathcal{L}_{d}$.
@@
 For the reconstruction network $\mathsf{M}^{r}$, we follow prior research in using Transformer encoder $\mathsf{T(\cdot)}$ for MMD [30, 28, 27, 29], formulated as:
-$[\mathbf{t}_{\text{CLS}},\mathbf{t}_{F}]=\mathsf{T}([\mathbf{CLS};\mathbf{F}]),\quad\mathbf{\hat{C}^{t}}=\mathbf{t}_{\text{CLS}}$ (6)
+$\mathbf{t}_{\text{CLS}},\mathbf{t}_{F}]=\mathsf{T}([\mathbf{CLS};\mathbf{F}]),\quad\mathbf{\hat{C}^{t}}=\mathbf{t}_{\text{CLS}}$ (6)
 where $CLS$ is a trainable classification token that serves as a global representation of all inputs, and its transformation is defined as the reconstructed caption embedding $\mathbf{\hat{C}^{t}}$.
@@
 1.
+Both input and target embeddings are L2-normalized, constraining representations to a unit hypersphere, allowing the MSE loss to focus on semantic alignment rather than magnitude variations.
 5.4.1 End-to-end (E2E) training
-To jointly optimize the reconstruction network $\mathsf{M}^{r}$ and the detection model $\mathsf{M}^{d}$, we explore end-to-end multi-task training, where the entire LAMAR model is trained using a joint loss function, which combines the reconstruction objective $\mathcal{L}_{r}$ (e.g., MSE) and the detection loss $\mathcal{L}_{d}$ (binary or categorical cross-entropy) as $\mathcal{L}=\mathcal{L}_{d}+\mathcal{L}_{r}$.
-This approach enables the detector to utilize both the input representation and the reconstructed embedding, while we explore various mechanisms to dynamically adjust the contribution of the reconstructed embedding based on relevance.
-Specifically, we explore:
+To jointly optimize the reconstruction network $\mathsf{M}^{r}$ and the detection model $\mathsf{M}^{d}$, we explore end-to-end multi-task training, where the entire LAMAR model is trained using both reconstruction loss $\mathcal{L}_{r}$ (e.g., MSE) and detection loss $\mathcal{L}_{d}$ (binary or categorical cross-entropy) as $\mathcal{L}=\mathcal{L}_{d}+\mathcal{L}_{r}$.
+Under this setting, the input caption embedding $\mathbf{C}$ may correspond to either a truthful or a manipulated caption, requiring $\mathsf{M}^{r}$ to learn how to handle both cases appropriately.
+Since $\mathsf{M}^{r}$ and $\mathsf{M}^{d}$ are trained jointly, the detection model can more dynamically modulate the influence of the reconstructed embedding based on its semantic alignment with the inputs.
+Specifically, we investigate:
 1.1) Direct integration of the reconstructed caption embedding $\mathbf{\hat{C}^{t}}$ into $\mathsf{M}^{d}$, as shown in Eq.5.
@@
 $\mathbf{\hat{C}^{t}}_{gate}=\mathbf{g}\odot\mathbf{\hat{C}^{t}}$ (8)
-where $\mathbf{g}$ is the gate, $S$ the sigmoid function and $\mathbf{W_{g}}\in\mathbb{R}^{768\times 768}$ while the input to Eq.5 is altered to $[\mathbf{F};\mathbf{\hat{C}^{t}}_{gate}]$.
+where $\mathbf{g}$ is the gate, $S$ the sigmoid function and $\mathbf{W_{g}}\in\mathbb{R}^{768\times 768}$ while the input to Eq.5 is altered to $\mathbf{F};\mathbf{\hat{C}^{t}}_{gate}]$.
 1.3) A masking mechanism, formulated as:
@@
 $\mathbf{\hat{C}^{t}}_{mask}=\mathbf{m}\odot\mathbf{\hat{C}^{t}}$ (10)
-where $\mathbf{m}$ represents binary mask sampled from a Bernoulli distribution, $\mathbf{W_{m}}\in\mathbb{R}^{768\times 768}$ and the input to Eq.5 is altered to $[\mathbf{F};\mathbf{\hat{C}^{t}}_{mask}]$.
+where $\mathbf{m}$ represents binary mask sampled from a Bernoulli distribution, $\mathbf{W_{m}}\in\mathbb{R}^{768\times 768}$ and the input to Eq.5 is altered to $\mathbf{F};\mathbf{\hat{C}^{t}}_{mask}]$.
 1.4) An attention mechanism, formulated as:
@@
 $\mathbf{\hat{C}^{t}}_{attend}=mean(softmax\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{T}}{\sqrt{768}}\right)\cdot\mathbf{V})$ (13)
-where $\mathsf{W}_{Q},\mathsf{W}_{K},\mathsf{W}_{V}\in\mathbb{R}^{768\times 768}$, and $mean$ denotes average pooling across the first dimension.
-The input to Eq.5 is altered to $[\mathbf{F};\mathbf{\hat{C}^{t}}_{attend}]$.
+where $\mathsf{W}_{Q},\mathsf{W}_{K},\mathsf{W}_{V}\in\mathbb{R}^{768\times 768}$, and meanmean denotes average pooling across the first dimension.
+The input to Eq.5 is altered to $\mathbf{F};\mathbf{\hat{C}^{t}}_{attend}]$.
 5.4.2 Large-scale Pre-Training (PT)
 In addition to end-to-end training, we also investigate a large-scale pre-training approach where the reconstruction network is trained exclusively on truthful captions using a large-scale image-caption dataset, VisualNews [20], comprising 1,259,732 truthful image-caption pairs.
+Unlike end-to-end training, in this setting $\mathsf{M}^{r}$ is trained separately using only perturbed captions $C^{f}$ as input.
 We explore two pre-training strategies:
@@
 For the “True vs.
-MC” experiments, we use the four versions of our “MisCaption This!” dataset, as detailed in Section 4, along with two additional datasets: the deduplicated version of the Crossmodal HArd Synthetic MisAlignment (CHASMA) dataset, which includes 145,891 truthful and 145,891 miscaptioned images [28], and the CLIP-based Named Entity Swapping by Topic (NESt) dataset, containing 847,693 miscaptioned images and 1,007,744 truthful images [17].
+MC” experiments, we use the 4 versions produced with LLaVA, 1 from Llama, 1 from Molmo, as detailed in Section 4, along with two additional datasets: the deduplicated version of the Crossmodal HArd Synthetic MisAlignment (CHASMA) dataset, which includes 145,891 truthful and 145,891 miscaptioned images [28], and the CLIP-based Named Entity Swapping by Topic (NESt) dataset, containing 847,693 miscaptioned images and 1,007,744 truthful images [30].
 For “True vs.
@@
 After training on any of the aforementioned synthetic datasets, we further assess performance on the VERITE benchmark [28], which comprises 1,000 real-world samples: 338 truthful image-caption pairs, 338 miscaptioned images, and 324 out-of-context pairs.
-For binary classification, we report “True vs.
-OOC” and “True vs.
-MC” accuracies, while for the multi-class task, we provide overall accuracy along with per-class accuracy.
+We report ‘True vs.
+OOC’ and ‘True vs.
+MC’ accuracies for binary classification tasks, and overall accuracy for the multi-class task.
 All three tasks serve as out-of-distribution evaluations, as the training data are synthetically generated, while the final evaluation is conducted on real-world data.
+For the ‘True vs.
+MC’ task, we update the VERITE benchmark, following the data selection criteria and labeling process defined in [28].
+Specifically, we collected 50 new samples – sourced from Snopes and Reuters, balanced across the two classes – describing events that occurred between October 2024 and July 2025.
+Crucially, these data points fall after the knowledge cut-off point of the selected VLMs; Molmo, Llama, and LLaVA.
+This expansion, which we denote VERITE 24/25, was collected to examine the out-of-distribution robustness and temporal generalization capacity of both zero-shot VLMs and our proposed methods against emerging misinformation.
 6.3 Competing Methods
-In addition to our proposed LAMAR architecture and its variations, we reproduce and evaluate the performance of several competing methods, without using any external evidence: (1) DT-Transformer [17]: a Transformer encoder that processes the input sequence $[\mathbf{CLS};\mathbf{I};\mathbf{C}]$; (2) RED-DOT (‘Baseline’ version) [21]: a Transformer-based model utilizing the fused representation $[\mathbf{CLS};\mathbf{F}]$ as input; (3) AITR (Attentive Intermediate Transformer Representations) [43], a model incorporating multimodal similarity (MUSE) and a self-attention mechanism over a stack of Transformer encoder blocks, with varying numbers of multi-head self-attention layers.
-We consider both its “attention pooling” and “weighted pooling” variants.
+In addition to our proposed LAMAR architecture and its variations, we reproduce and evaluate the performance of several competing methods: (1) MUSE [29]: a simple similarity-based baseline consisting of a single hidden layer applied to cosine similarity scores between image-text pairs.
+(2) SpotFake* [34]: A neural network that processes the input sequence $\mathbf{I};\mathbf{C}]$.
+(3) SAFE* (Similarity-Aware FakE) [43]: Similar to SpotFake*, but processing the input sequence $\mathbf{I};\mathbf{C};sim(\mathbf{I},\mathbf{C})]$.
+(4) MVAE* (Multimodal Variational Autoencoder) [17]: A bimodal variational autoencoder that maps image and text features into a shared latent space.
+The network is trained to jointly minimize the reconstruction, KL divergence, and cross-entropy losses.
+The decoder reconstructs the original input words and the image embeddings from the shared latent space.
+(5) BCMF* (Bidirectional Cross-Modal Fusion) [40]: A dual-stream architecture that performs bidirectional cross-attention between images and captions, followed by gating mechanisms to refine the features.
+The final fused feature vector concatenates the refined features and their explicit interaction terms, expressed as $\text{Maxpool}(\mathbf{C}^{\prime});\mathbf{I}^{\prime};|\text{Maxpool}(\mathbf{C}^{\prime})-\mathbf{I}^{\prime}|]$, which is then used for the final classification.
+(6) PMFN* (Progressive MLP-Mixer Fusion Network) [16]: A multi-stage architecture, that performs hierarchical fusion.
+It leverages the MLP-Mixer module to iteratively refine multimodal features by fusing image and text representations from different abstraction stages of the backbones, aiming to enhance the correlation between shallow and deep features before final classification.
+(7) DT-Transformer [30]: a Transformer encoder that processes the input sequence $\mathbf{CLS};\mathbf{I};\mathbf{C}]$.
+(8) RED-DOT [27]: a Transformer-based model operating on the fused representation $\mathbf{CLS};\mathbf{F}]$; we use the Baseline version without external evidence.
+(9) AITR (Attentive Intermediate Transformer Representations) [29]: A model that uses a stack of $n$ Transformer encoder layers of varying number of attention heads to process the initial multimodal sequence $\mathbf{CLS}_{0};\mathbf{F}]$ and then applies a self-attention mechanism over the sequence of intermediate classification tokens $\mathbf{CLS}_{1},\dots,\mathbf{CLS}_{n}]$ for final classification.
+For a fair comparison, all models marked with an asterisk (*) were adapted to use CLIP ViT-L/14 image and text features, replacing their original backbones (VGG-19, BERT, DeiT, Swin Transformer, Bi-LSTM, etc.).
 6.4 Implementation Details
 We train LAMAR using a Transformer encoder with 4 encoder blocks, each containing 4 multi-head self-attention heads and a feed-forward layer of 1,024 dimensions and a dropout of 0.1 probability.
-The model is optimized with Adam, using a learning rate of $41e-$ 4 and a batch size of 512.
+The model is optimized with Adam, using a learning rate of 1e41e-4 and a batch size of 512.
 We train the network for up to 50 epochs, with early stopping after 10 epochs if validation performance does not improve.
 For the large-scale pre-training, we consider $\sigma\in\{0.1,0.2\}$ and $\mu=0.0$ and dropout probability $dp\in\{0.2,0.5\}$.
-To ensure reproducibility, we set a constant random seed of 0 for PyTorch, Python Random, and NumPy.
+To ensure reproducibility, we set a constant random seed (0) for PyTorch, Python Random, and NumPy.
 6.5 Computational Complexity
-Generating a single version of “MisCaption This!”  including the generation of captions for 35,536 images, took 18.3 hours using a single Nvidia GeForce RTX 4090 (24GB RAM).
-Feature extraction with CLIP ViT-L/14 required approximately 16 minutes for images and 2 minutes for texts, using a batch size of 256.
-Both data generation and feature extraction were repeated four times for each version of “MisCaption This!”.
+Generating a “MisCaption This!” dataset with LLaVA, took 18.3 hours using a single Nvidia GeForce RTX 4090 (24GB RAM).
+Feature extraction with CLIP ViT-L/14 required approximately 16 minutes for images and 2 minutes for texts, using a batch size of 256; and was performance only once per dataset, and is used for any model.
 Using fvcore [^5] https://github.com/facebookresearch/fvcore, we estimate that LAMAR, trained end-to-end with a gate mechanism, requires 980,997 FLOPs, and 1,034,181 FLOPs with the attention mechanism.
-This is comparable to AITR (1,070,653 FLOPs) [43] and significantly more efficient than RED-DOT (2,208,072 FLOPs) [21].
-When using pre-extracted features, LAMAR requires a maximum of 30 seconds per epoch (batch size 512) to process the full “MisCaption This!” for multiclass classification.
-Figure 4: Performance of detection models (DT-Transformer and RED-DOT) trained on four variations of the “MisCaption This!” dataset ($\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\mathcal{D}_{3}$, and $\mathcal{D}_{4}$), evaluated under varying filtering thresholds ($l\in{0,5,10,15,25,50,\text{None}}$) in terms of test-set accuracy and VERITE “True vs.
+This is comparable to AITR (1,070,653 FLOPs) [29] and more efficient than RED-DOT (2,208,072 FLOPs) [27].
+When using pre-extracted features, LAMAR requires a maximum of 30 seconds per epoch (batch size 512) to process the unfiltered ($l=None$) dataset for multiclass classification.
+7 Results
+Figure 5: Performance of detection models (DT-Transformer and RED-DOT) trained on four variations of LLaVA-$\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\mathcal{D}_{3}$, and $\mathcal{D}_{4}$, evaluated under varying filtering thresholds ($l\in\{0,5,10,15,25,50,\text{None}\}$, or 4.5%, 19.1%, 27.8%, 34.9%, 47.9%, 74.0%, and 100% of the dataset) in terms of Test-set Accuracy and VERITE “True vs.
 MC” Accuracy.
-7 Results
 7.1 Dataset Variants and Filtering
-We first examine the generalizability of two established MMD models: DT-Transformer [17] and RED-DOT [21] (without incorporating external evidence) when trained on the four “MisCaption This!” variations ($\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\mathcal{D}_{3}$, and $\mathcal{D}_{4}$), under different filtering thresholds $l\in\{0,5,10,15,25,50,\text{None}\}$.
+First, we investigate how different synthetic data variations ($\mathcal{D}_{1}$$\mathcal{D}_{4}$) and filtering thresholds ($l$) influence a model’s ability to detect synthetic and real-world misinformation.
+Specifically, we analyze the relationship between internal (synthetic) test-set performance and out-of-distribution generalization to the VERITE (“True vs.
+MC”) benchmark.
 As shown in Fig.
-4, test-set accuracy tends to improve with higher $l$ values across all datasets and models.
-The highest performance is observed with no filtering ($l=\text{None}$), ranging from 79% to 83.4%.
-In contrast, evaluating models on the real-world data of VERITE (“True vs.
-MC”), there is a slight negative (Pearson) correlation between test-set accuracy and VERITE accuracy: -0.29 for DT-Transformer and -0.17 for RED-DOT.
-This suggests that higher test-set accuracy does not guarantee better generalization, as training with synthetic misinformation that is too easily distinguishable from truthful pairs may lead to inflated test-set accuracy without improving performance on real-world misinformation.
-Both models achieve peak performance at $l=10$, with accuracies of 63.3% and 61.8%, followed by $l=5,15,$ and $25$; highlighting the importance of filtering out “rambling” and generic descriptions sometimes produced by LLaVa.
+5, test-set accuracy tends to improve with higher $l$ values across all datasets and two established MMD models, DT-Transformer [30] and RED-DOT [27].
+The highest test-set performance is observed without any filtering ($l=\text{None}$), ranging from 79% to 83.4%.
+However, we identify a divergence between test-set and real-world performance: with a negative Pearson correlation of 0.29-0.29 for DT-Transformer and 0.17-0.17 for RED-DOT.
+This indicates that high test-set accuracy on synthetic data does not guarantee better generalization; rather, training on synthetic misinformation that is too easily distinguishable from truthful pairs likely encourages the model to exploit spurious shortcuts rather than learning robust factual consistency, ultimately hindering its generalization.
+Regarding filtering, both models achieve peak performance at $l=10$, with accuracies of 63.3% and 61.8%, followed by $l=5$, $15$, and $25$.
+This highlights the importance of filtering out the “rambling” or generic descriptions occasionally produced by VLMs.
+Moreover, filtering overly long texts prevents the detection model from relying on superficial length-based patterns (e.g., that longer texts tend to be “falsified”) rather than actual factual consistency.
 Thus, by refining the dataset in this way, we improve model robustness and enhance real-world generalization on VERITE.
 Overall, both models achieve their highest performance when trained on $\mathcal{D}_{3}$, with RED-DOT and DT-Transformer averaging 60.2% and 59.2%, respectively.
-This suggests that $p_{3}^{gen}$ generates misinformation patterns that more closely resemble real-world examples, leading to better model generalization.
-Therefore, for subsequent experiments, we will continue using the $\mathcal{D}_{3}$ version of “MisCaption This!”, while treating $l$ as a tunable hyper-parameter.
-Table 1: Comparative and ablation analysis of the proposed LAMAR method and prior SotA models (DT-Transformer, RED-DOT, and AITR) trained on three synthetic datasets (“MisCaption This!” ($\mathcal{D}_{3}$), NESt, and CHASMA) and evaluated on VERITE (“True vs.
-MC”).
-Bold denotes the highest accuracy on VERITE.
-Model $\mathcal{D}_{3}$ NESt CHASMA
-DT-Transformer 61.8 58.7 57.8
-RED-DOT 63.3 57.0 58.0
-AITR 62.9 58.4 58.3
-LAMAR Variant $\mathcal{D}_{3}$ NESt CHASMA
+This suggests that $p_{3}^{gen}$ produces misinformation patterns that more closely resemble real-world examples, resulting in better model generalization.
+Consequently, we adopt $p_{3}^{gen}$ to produce Molmo-$\mathcal{D}_{3}$ and Llama-$\mathcal{D}_{3}$, and use LLaVA-$\mathcal{D}_{3}$ variant of “MisCaption This!” for proceeding experiments, while treating $l$ as a hyperparameter.
+7.2 Ablation across Datasets
+Table 1: LAMAR ablations trained on three synthetic datasets (LLaVA-$\mathcal{D}_{3}$, NESt, or CHASMA) and evaluated on VERITE (“True vs. MC”). Bold denotes the highest accuracy.
+LAMAR Variant LLaVA-$\mathcal{D}_{3}$ NESt CHASMA
 E2E, Gate 66.0 61.2 59.8
@@
 E2E, Mask 63.6 59.5 58.1
+E2E, Gate - Fusion 61.1 57.3 58.0
+E2E, Attention - Fusion 61.2 55.0 57.0
 E2E, Direct 63.8 59.2 57.8
@@
 E2E, Direct - Text 63.0 58.7 57.1
-No Reconstruction 62.7 58.2 56.7
 PT, Dropout, Gate 63.0 57.5 58.6
@@
 PT, Gaussian, Direct 63.2 58.6 57.5
+No Reconstruction 62.7 58.2 56.7
+We conduct an extensive analysis of the proposed LAMAR method, including various model variants and ablations.
+This analysis is conducted using three distinct synthetic datasets for training: “MisCaption This!” (LLaVA-$\mathcal{D}_{3}$), CHASMA [28], and NESt [30].
+The performance of all models is consistently evaluated on the VERITE benchmark, specifically, the “True vs.
+MC” task.
+As shown in Table 1, training on “MisCaption This!” (LLaVA-$\mathcal{D}_{3}$) consistently yields models with stronger generalization to real-world data.
+This trend holds across all LAMAR variants and supports our hypothesis that VLM-generated training data provides more effective supervision than earlier methods based on named entity swaps or cross-modal inconsistencies.
+The best overall performance (66%) is achieved by LAMAR [E2E] with the Gate integration mechanism trained on “MisCaption This!” data (LLaVA-$\mathcal{D}_{3}$), outperforming NESt by 7.8% and CHASMA by 10.4%.
+E2E training proves advantageous, as it directly optimizes reconstruction over actual miscaptioned pairs, unlike PT variants that rely on perturbed truthful captions via noise or dropout.
+These results highlight the value of semantically grounded, VLM-generated training data that more closely resembles real-world misinformation.
+Among integration strategies, Gate and Attention mechanisms consistently outperform direct embedding injection, highlighting the importance of adaptive integration.
+In contrast, the Mask mechanism underperforms, likely due to reduced representational capacity from fully masking the reconstruction embedding.
+Similarly, removing the fusion component $\mathbf{F}$ (Eq.
+4) and using simple concatenation ($\mathbf{I};\mathbf{C}]$) significantly degrades performance, emphasizing the role of effective multimodal fusion.
+We also conduct unimodal ablations.
+Removing the image input (Direct-Image) leads to a slight drop in performance (-0.9%), while removing the text (Direct-Text) results in a larger drop (-1.3%), effectively turning the task into latent captioning.
+Finally, eliminating the reconstruction module entirely yields the lowest E2E performance, confirming the critical role of the reconstruction in guiding accurate detection.
+To evaluate how LAMAR integrates the reconstructed signal with the primary inputs, we analyze the distribution of attention scores across input tokens (image, caption, and reconstructed caption embeddings) and the behavior of the gating mechanism.
+Attention scores are substantially higher on the reconstruction token compared to the image or input caption tokens across classes.
+Truthful pairs receive higher attention on average for both the image token (0.218 vs.
+0.176) and the input caption (0.213 vs.
+0.169), while falsified pairs shift attention toward the reconstructed caption (0.655 vs.
+0.569).
+Similarly, the gate score correlates positively with reconstruction loss (0.451), and reconstruction loss itself is positively associated with the model’s predicted label (0.386), suggesting that higher reconstruction error pushes the model toward a “falsified” prediction.
+These patterns indicate that LAMAR learns to focus more on the reconstructed caption, via attention or gating, when detecting inconsistencies, making it a valuable auxiliary signal for accurate classification.
+Thus, when the reconstruction caption embeddings diverges semantically from the input embeddings, the model flags the pair as falsified.
+7.3 Comparative Study across VLM-Generated Datasets
+Table 2 shows the performance of LAMAR and prior SotA models trained on VLM-generated datasets produced by LLaVA, Molmo, and Llama, and evaluated on the VERITE benchmark (“True vs.
+MC”).
+Across all three datasets, LAMAR consistently outperforms prior models by significant margins: 14.4% over MUSE, 8.7% over MVAE, 7.0% over SpotFake, 6.3% over SAFE, 3.8% over BCMF, 3.8% over MPFN, 6.8% over DT-Transformer, 4.3% over RED-DOT, and 4.9% over AITR.
+This holds for both LAMAR variants using gated or attention-based integration.
+The best overall result is achieved by LAMAR [E2E, Gate] trained on LLaVA-$\mathcal{D}_{3}$, reaching 66.0% accuracy.
+Table 2: Performance of models trained on three VLM-generated datasets (LLaVA-$\mathcal{D}_{3}$, Molmo-$\mathcal{D}_{3}$, Llama-$\mathcal{D}_{3}$) and evaluated on VERITE (“True vs.
+MC”).
+Bold indicates the highest overall accuracy; Underline indicates the best accuracy per training dataset.
+Model LLaVA-$\mathcal{D}_{3}$ Molmo-$\mathcal{D}_{3}$ Llama-$\mathcal{D}_{3}$
+MUSE [29] 57.7 54.9 57.5
+MVAE [17] 60.7 60.8 60.5
+SpotFake [34] 61.7 62.6 60.7
+SAFE [43] 62.1 63.0 59.6
+BCMF [40] 63.6 61.2 59.0
+MPFN [16] 63.6 64.1 58.6
+DT-Transformer [30] 61.8 62.1 58.4
+RED-DOT [27] 63.3 62.9 61.7
+AITR [29] 62.9 63.2 63.0
+LAMAR [E2E, Attention] 65.1 65.1 64.2
+LAMAR [E2E, Gate] 66.0 64.4 63.9
+LLaVA-generated data yielding higher performance may be partially attributed to the use of “adversarial prompt selection” on it, whereas Molmo and Llama data were generated using the same prompt ($p_{3}^{gen}$).
+While this introduces a slight advantage for LLaVA, it is notable that models trained on Molmo- and Llama-generated data performed comparably well, thus underscoring the general utility of VLM-generated data across models.
+Qualitative analysis (see Fig.
+3) further highlights that LLaVA-$\mathcal{D}_{3}$ contain more nuanced and believable false captions by maintaining semantic consistency with the image.
+In contrast, LLaVA-$\mathcal{D}_{4}$ and Llama-$\mathcal{D}_{3}$ often contain more apparently false captions – for example, attributing protests to “Madrid, Spain” or “in front of the Eiffel Tower in Paris” despite clear visual cues such as Greek flags.
+Such easily detectable falsehoods are less reflective of real-world misinformation and thus less effective for training models that generalize well.
+7.4 Out-of-context and Multiclass Detection
 Table 3: Performance of models on the “True vs. OOC” task, trained on the NewsCLIPpings dataset and evaluated on both NewsCLIPpings and VERITE.
@@
 SSDL [23] 71.0 -
-DT-Transformer 79.7 69.4
-RED-DOT 81.5 73.5
-AITR 84.1 74.1
+MVAE [17] 60.6 55.2
+MUSE [29] 80.6 72.0
+SpotFake [22] 79.8 67.1
+SAFE [43] 83.7 72.7
+BCMF [40] 76.6 67.1
+MPFN [16] 76.1 66.5
+DT-Transformer [30] 79.7 69.4
+RED-DOT [27] 81.5 73.5
+AITR [29] 84.1 74.1
 LAMAR [E2E, Attention] 84.8 75.1
 LAMAR [E2E, Gate] 84.7 76.3
-7.2 Ablation across Datasets
-We train three established MMD models, DT-Transformer, RED-DOT, and AITR, alongside various variants and ablations of the proposed LAMAR method using three synthetic “True vs.
-MC” training datasets: “MisCaption This!” ($\mathcal{D}_{3}$), CHASMA [28], and NESt [17], and evaluating their performance on VERITE (“True vs.
-MC”).
-As shown in Table 1, using “MisCaption This!” ($\mathcal{D}_{3}$) as the training dataset consistently results in models with higher generalizability to real-world data (VERITE).
-This trend holds across all evaluated models and variants.
-These results support our hypothesis that LVLMs are capable of generating more effective synthetic training data, ultimately leading to more robust detection models, compared to previous approaches that rely on named entity swaps or cross-modal misalignment.
-Furthermore, we observe that LAMAR trained end-to-end (E2E) with the “Gate” mechanism for integration achieves the highest overall performance on VERITE across all three datasets, yielding the best performance (66%) when using the “MisCaption This!” ($\mathcal{D}_{3}$); achieving 7.8% higher performance compared to when trained on NESt, and 10.4% higher compared to CHASMA.
-Overall, we observe that E2E training tends to yield better performance, as it is directly optimized to reconstruct the embeddings of actual miscaptioned images, in contrast to PT methods that reconstruct embeddings of truthful pairs interjected with Gaussian noise or dropout.
-Additionally, LAMAR with E2E training, utilizing either the “Gate” or “Attention” mechanisms for integrating the reconstruction embeddings, consistently outperforms the direct integration; emphasizing the importance of utilizing an integration mechanism.
-The “Mask” mechanism does not perform as well, likely due to the complete masking of the reconstruction embedding, which reduces its expressiveness.
-Instead, modulating the embedding by adjusting its values based on learned gate values or attention scores proves to be a more effective strategies.
-We also investigate unimodal reconstruction as an ablation.
-Direct integration without the image input (Direct-Image), using only the text as input, results in a slight decrease in performance (-0.9%).
-Similarly, excluding textual information (Direct-Text), which effectively transforms the task into a latent image captioning approach, leads to even lower performance (-1.3%).
-Removing the reconstruction network entirely from LAMAR, leaving only the detection classifier, yields the lowest performance among all end-to-end (E2E) approaches.
-Finally, we observe that LAMAR [E2E, Gate] outperforms prior state-of-the-art models, such as DT-Transformer, RED-DOT, and AITR, by 6.8%, 4.3%, and 4.9%, respectively, when trained on “MisCaption This!” ($\mathcal{D}_{3}$); validating the effectiveness of our proposed method for the detection of miscaptioned images.
-Table 3: Performance of models trained on the multi-class “MisCaption This!” ($\mathcal{D}_{3}$) or the combined CHASMA and NewsCLIPpings datasets, evaluated on VERITE, reported as overall accuracy and per-class accuracy.
-Training Model Accuracy True MC OOC
-CHASMA DT-Transformer 50.0 78.7 23.1 48.0
-RED-DOT 48.5 71.9 21.9 51.7
-AITR 51.4 90.5 15.4 48.0
-LAMAR [E2E, Gate] 53.2 89.6 18.1 51.7
-$\mathcal{D}_{3}$ DT-Transformer 47.8 53.0 60.4 29.2
-RED-DOT 48.8 50.9 53.3 41.9
-AITR 51.7 50.3 62.7 41.5
-LAMAR [E2E, Gate] 54.2 58.6 58.6 44.9
-7.3 Out-of-context and Multiclass Detection
-Table 2 presents the performance of LAMAR compared to prior SotA models on the binary classification task of “True vs.
-OOC”, without external evidence.
-LAMAR achieves the highest accuracy on the NewsCLIPpings dataset, with LAMAR [E2E, Attention] reaching 84.8% and LAMAR [E2E, Gate] closely following at 84.7%, outperforming AITR (84.1%) and other models.
-More importantly, when evaluated on VERITE, LAMAR [E2E, Gate] attains 76.3%, surpassing the best prior model (AITR) by 3%.
-We observe a noticeable performance gap between “True vs.
-OOC" (76.3%) and “True vs.
-MC" (66.0%) due to differing task complexity.
-In OOC, the entire image is mismatched with the caption, making discrepancies more apparent, while MC involves subtle manipulations (e.g., actions, dates, locations, names), making it harder to distinguish from truthful captions.
-For the multi-class classification task, we consider two training datasets: “MisCaption This!” ($\mathcal{D}_{3}$) and CHASMA combined with NewsCLIPpings.
-As shown in Table 3, LAMAR [E2E, Gate] continues to demonstrate superior generalization, regardless of the dataset used for training.
-When trained on “MisCaption This!” ($\mathcal{D}_{3}$), LAMAR achieves the highest overall accuracy (54.2%), representing a relative improvement of 4.8%, 11.1%, and 13.4% over AITR, RED-DOT, and DT-Transformer, respectively.
-Notably, all four models achieve relatively higher accuracy on the MC class when trained on “MisCaption This!” ($\mathcal{D}_{3}$) compared to CHASMA+NewsCLIPpings.
-These results validate the effectiveness of our proposed method in leveraging synthetic data generated by an LVLM and highlight the impact of incorporating a reconstruction network to enhance the detection process, ultimately improving generalization to real-world misinformation detection.
-8 Conclusions
-In this study, we address the challenge of MMD by introducing the “MisCaption This!” dataset, and the Latent Multimodal Reconstruction (LAMAR) method.
-To create “MisCaption This!”, we use an LVLM to generate diverse and realistic synthetic miscaptioned images, enhancing training data quality and by extension model generalization.
-LAMAR employs a reconstruction-based approach, where the original truthful caption embeddings are reconstructed from manipulated image-caption pairs, providing an auxiliary signal to aid the detection process.
-Extensive experiments show that models trained on “MisCaption This!” achieve superior out-of-distribution generalization on the real-world VERITE benchmark, outperforming named entity manipulations and cross-modal misalignment by 7.8% and 10.4%, respectively.
-Additionally, LAMAR set new state-of-the-art performance on all three VERITE tasks—“True vs.
-OOC”, “True vs.
-MC”, and multiclass classification—outperforming prior SotA by 4.3%, 3.0%, and 4.8%, respectively.
-These results highlight the potential usefulness of leveraging LVLMs for dataset generation and the potential of reconstruction networks in improving MMD performance.
-Our study represents a first step toward leveraging LVLMs to enhance training data quality for MMD but is limited to leveraging a single model, LLaVa 1.6 7B.
-Future research could explore alternative LVLMs, jailbreak prompts, prompt tuning or few-shot prompting to further improve the robustness of generated training data.
-Nevertheless, researchers should consider potential risks, particularly in inadvertently enabling malicious actors to refine AI-generated disinformation campaigns.
-Furthermore, future research could further refine the reconstruction network by integrating additional external information or evidence to enhance the detection process.
-We do not explore this avenue due to the absence of datasets containing external evidence for miscaptioned images.
-A key challenge in collecting and leveraging external information is the risk of information leakage, which can undermine the early detection of emerging misinformation [42, 31].
-Additionally, existing evidence-based OOC datasets often contain artifacts that models may exploit as shortcuts rather than truly assessing factual consistency [43].
-Addressing these limitations would be a valuable step forward for the fields of MMD and automated fact-checking.
+Table 4: Comparison of multi-class models trained either on CHASMA, LLaVA-$\mathcal{D}_{3}$, Molmo-$\mathcal{D}_{3}$, and evaluated on VERITE.
+Model CHASMA LLaVA-$\mathcal{D}_{3}$ Molmo-$\mathcal{D}_{3}$
+MUSE [29] 38.5 49.7 47.3
+MVAE [17] 39.8 40.5 42.7
+SpotFake [34] 49.2 48.7 47.2
+SAFE [43] 50.4 51.4 51.0
+BCMF [40] 45.9 45.4 47.5
+MPFN [16] 46.1 44.1 44.8
+DT-Transformer [30] 50.0 47.8 48.7
+RED-DOT [27] 48.5 48.8 49.1
+AITR [29] 51.4 51.7 52.0
+LAMAR [E2E, Gate] 53.2 54.2 54.9
+Table 3 presents LAMAR’s performance compared to prior SotA models on “True vs.
+OOC” classification, without access to external evidence.
+LAMAR achieves the highest accuracy on the NewsCLIPpings dataset, with [E2E, Attention] reaching 84.8% and [E2E, Gate] closely following at 84.7%, both outperforming AITR (84.1%) and other models.
+More importantly, LAMAR [E2E, Gate] achieves 76.3% on VERITE, surpassing the best prior model (AITR) by 3%.
+We observe a consistent performance gap between True vs.
+OOC” (76.3%) and True vs.
+MC” (66.0%), reflecting the greater difficulty of detecting subtle manipulations (e.g., dates, places, or actions) in MC cases, compared to the more blatant mismatches typical of OOC examples.
+For the multi-class classification task, we employ three training datasets: LLaVA-$\mathcal{D}_{3}$, Molmo-$\mathcal{D}_{3}$, and CHASMA; all being combined with NewsCLIPpings to represent the OOC class.
+As shown in Table 4, LAMAR [E2E, Gate] consistently generalizes best across all training sets.
+When trained on Molmo-$\mathcal{D}_{3}$, LAMAR achieves the highest overall accuracy (54.9%), representing a relative improvement of 5.6% over AITR.
+These results further validate the benefits of using VLM-generated data in LAMAR.
+Table 5: Comparison of zero-shot VLMs against fine-tuned LAMAR variants on the “True vs. MC” task of VERITE and the post-knowledge-cutoff data of VERITE 24/25.
+Model Parameters VERITE VERITE 24/25
+Llama 3.2 11B 59.3 53.1
+Molmo 7B 63.0 58.0
+LLaVA 1.6 7B 66.9 62.0
+LAMAR [E2E, Attention] 21M 65.1 70.0
+LAMAR [E2E, Gate] 19M 66.0 68.0
+7.5 Temporal Generalization Analysis
+Table 5 compares the performance of the three VLMs used for the generation of “MisCaption This!” data, in a zero-shot setting against our fine-tuned LAMAR variants on the VERITE “True vs.
+MC” task.
+On the original VERITE dataset, the performance of the VLMs varies, with LLaVA 1.6 (66.9%) showing comparable performance to LAMAR [E2E Gate] (66.0%) and outperforming Molmo (63.0%) and Llama 3.2 (59.3%).
+Decent zero-shot performance is not unexpected, as these VLMs are trained on vast, large-scale datasets, making it highly probable that they have encountered samples or related content from popular fact-checking sources like Snopes and Reuters used in the original VERITE benchmark.
+However, the analysis shifts when evaluated on the VERITE 24/25 set comprising samples dating after the VLMs’ knowledge cut-off.
+Here, all zero-shot VLMs demonstrate a notable decline in performance, with LLaVA 1.6 dropping to 62.0%, Llama 3.2 falling to 53.1%, and Molmo to 58.0%.
+In contrast, both fine-tuned LAMAR variants demonstrate strong temporal generalization, with LAMAR [E2E Attention] achieving 70.0% and LAMAR [E2E Gated] 68.0% on the post-cutoff data, despite having only $\approx$20 million parameters compared to the 7-11 billion parameters of the VLMs.
+The performance gap on VERITE 24/25 highlights that VLM’s zero-shot performance is limited by their training data and knowledge cutoff, whereas LAMAR’s reconstruction-based approach provided higher robustness against emerging misinformation.
+8 Ethical Discussion
+Our work demonstrates the potential of VLMs for generating high quality training datasets for multimodal misinformation detection.
+Nevertheless, it is crucial to consider the ethical implications of this work and future research in this direction.
+In our initial attempts, earlier open-source VLMs such as MiniGPT and LLaVA 1.5 struggled to generate coherent or realistic false captions.
+However, later versions (LLaVA 1.6, Llama 3.2, and Molmo) were more capable of producing false but plausible captions.
+While this capability proved valuable for research – enabling the creation of more robust training datasets – it also raises serious societal concerns as even lightweight, open-source models can now be repurposed to automate and scale disinformation campaigns.
+We believe that the research community must remain vigilant, ensuring that advancements in detection do not inadvertently provide a blueprint for malicious generation.
+To mitigate these risks, we refrain from publicly releasing specific prompts used to generate the “MisCaption This!” data.
+However, to ensure scientific reproducibility, we provide the general templates within this work, and the exact prompts will be made available to verified researchers upon reasonable request for non-commercial research purposes.
+9 Conclusion
+In this study, we address the challenge of Multimodal Misinformation Detection (MMD) by introducing the “MisCaption This!” framework and the Latent Multimodal Reconstruction (LAMAR) architecture.
+“MisCaption This!” leverages Vision-Language Models (VLMs) to generate diverse, realistic synthetic miscaptioned images, significantly improving training data quality and ultimately model generalization.
+The LAMAR architecture adopts a reconstruction-based approach: it learns to reconstruct the embedding of the original, truthful caption from manipulated image-caption pairs, thereby offering an effective auxiliary signal that enhances detection.
+Extensive experiments demonstrate that models trained on “MisCaption This!” data significantly outperform previous methods – including those based on named entity manipulation and cross-modal misalignment – by 7.8% and 10.4% on the VERITE benchmark, respectively.
+Furthermore, LAMAR establishes new SotA performance across all three VERITE tasks: True vs.
+OOC (+4.3%), True vs.
+MC (+3.0%), and multiclass classification (+5.6%).
+Despite the promising gains achieved by “MisCaption This!” and LAMAR, this work stresses that MMD remains an open and challenging task.
+Apart from high “out-of-context” performance ($\approx 85\%$), our final model achieved $\approx 66\%$ and $\approx 55\%$ in the miscaptioned and multiclass scenarios, respectively, which indicates considerable room for improvement.
+We argue that while assessing the internal consistency of the image-caption pair is a crucial prerequisite for MMD, it is rarely sufficient.
+Therefore, a critical future direction involves collecting and integrating external knowledge and evidence (e.g., related news articles) for comprehensive fact-checking.
+Although evidence-enriched datasets exist, such as the enriched NewsCLIPpings dataset for out-of-context cases [1], they often suffer from significant issues including information leakage, dataset-specific biases [29], or unreliable evidence [10, 9].
+This emphasizes the pressing need for a new, robust, evidence-enriched dataset specifically targeting miscaptioned images to accurately model and address external inconsistency.
+Synthetic data generation through “MisCaption This!”, and reconstruction-based architectures, such as LAMAR, could play an instrumental role in advancing these research directions.
 Acknowledgments
-This work is partially funded by the project “vera.ai: VERification Assisted by Artificial Intelligence” under grant agreement no.
-101070093 and “DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies” under grant agreement no.
-101079164.
+This work is partially funded by the projects “vera.ai: VERification Assisted by Artificial Intelligence” under grant agreement no.
+101070093, “DisAI - Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies” under grant agreement no.
+101079164, and “AI4TRUST - AI-based-technologies for trustworthy solutions against disinformation” under grant agreement no.
+101070190.